### v2

# Price Experimentation and Interference

[…]
Price Experimentation and Interference
[…]
In this paper, we examine biases arising in A/B tests where firms modify a continuous parameter, such as price, to estimate the global treatment effect of a given performance metric, such as profit. These biases emerge in canonical experimental estimators due to interference among market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to a standard revenue management pricing problem. This setting highlights a key pitfall in the use of A/B pricing experiments to guide profit maximization: notably, the canonical estimator for the expected change in profits can have the wrong sign. In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo. We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup.Finally, we discuss structural and practical implications for platform operators.
[…]
1 Introduction
In this work we analyze the biases that arise when a technological firm or platform conducts A/B tests by varying a continuous parameter to determine the global treatment effect associated to a performance metric of interest. These biases appear when using canonical designs and estimators and result from interference between market participants.
More concretely, consider the example of a canonical pricing experiment where a fraction of the products are assigned a price that is higher than the baseline price. The global treatment effect of this higher price compares two counterfactual universes: one where all products have the high price (global treatment), and one where all products have the low price (global control). Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price; when using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect. This phenomenon is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts; see Related Work (Section 2) for background.
Our paper considers the following question: Given the standard experimental designs (e.g., product-level or customer-level randomization) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments? There are several key features of our approach. First, drawing inspiration from recent work in the operations literature, we employ structural modeling to gain valuable insights into biases. Furthermore, our analysis focuses on differential changes in continuous variables such as prices, enabling us to leverage calculus to derive straightforward and intuitive expressions. (In practical settings, we will think of “small” changes around the control value.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
We first introduce a succinct yet general model of platform experimentation in which a firm wants to understand the effect of a continuous variable on a metric of interest. The model is general in that can it encode different market arrangements and experimental designs. By leveraging differential calculus, we provide several structural results regarding the performance of canonical estimators and designs. Notably, we show that the bias corresponds to the cross-effects of the control and treatment metrics with respect to marginal changes on the other group’s variable. These cross-effects are precisely an indicator of interference, providing an intuitive structural characterization of bias.
Second, we specialize our results to a canonical revenue management pricing problem. By using our bias characterizations, we show that experiments can obtain the wrong sign when the metric of interest is profit (or revenue). In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive). (Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate obtains the correct sign.) We express the conditions under which an estimate with incorrect sign can be obtained with straightforward expressions that involve price elasticities. [^1] The expressions are reminiscent of elasticity conditions in Blake and Coey, 2014a . Our results have important implications for price optimization based on experiments: incorrectly following the experiment’s suggestion may lead one to move in the wrong direction and decrease profits relative to the status-quo!
We apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs. For example, for a listing-side experiment a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low. We also calibrate this stylized model to experimental estimates from Airbnb (published in Holtz et al., 2020), and investigate empirical implications for regimes where a change of sign may occur in a price experiment.
2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, farias2023correcting), and message passing (shirani2023causal).
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023). Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). roemheld2023interference uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper – those where a continuous variable (such as price) is incrementally varied – as marginal policy changes, and develops a framework to study them.
The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “naive” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments. In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price in our present paper. Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is more focused on design and estimation for general estimands in market equilibrium. (An earlier version of Munro et al., (2021) considers treatment changes in a continuous variable, with an emphasis on relating the resulting differential treatment effect to direct and indirect effects.)
3 General Model Formulation
This section presents our general model, which we subsequently adapt to specific settings.
3.1 Market Model and Estimand
We consider a market in which a firm or platform sets a single variable of interest, $x\in\mathbb{R}_{+}$, evaluated by a performance metric $T(x)\in\mathbb{R},\forall x\in\mathbb{R}_{+}$. The performance metric can describe bookings, profits, or revenues of the platform. We make the following assumption that is kept throughout the paper. {assumption} The function $T(x)$ is differentiable for all $x\in\mathbb{R}_{+}$.
The firm is interested in optimizing the performance metric by adjusting the decision variable $x$. In particular, the firm is interested in the difference in the metric when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, relative to a status-quo “control” value, denoted by $x_{0}$. This difference, $T(x_{1})-T(x_{0})$, corresponds to the global treatment effect ($GTE$). We define a local version of $GTE$ considering that the treatment value of the decision variable is “close” to the control value:
$GTE(x_{0})$ ::=:= $T^{\prime}(x_{0})\ .$ (1)
$GTE(x_{0})$ is our estimand.
3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be modified to optimize the performance metric. We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated in each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
It is worth noting that in the settings we are considering, randomization can occur at either side of the market, i.e., customers or products (i.e., listings in the case of a platform) can be randomized. In a customer-side experiment, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. On the other hand, in a product (listing)-side experiment, a customer is exposed to a mix of treatment and control products (or listings). As before, we denote by $x_{0}$ the baseline (control) value of the decision variable and by $x_{1}$ the treatment value.
We assume that the performance metric under the treatment and control conditions are given by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$, which we refer to as the control and treatment performance metrics, respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment value are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units is randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
The canonical experimental designs at each side of the market have an associated natural “naive” local estimator. To define this estimator, consider a sequence of experiments for which the treatment value of the decision variable approaches the control value. Then, we compute the relative difference between the performance metrics appropriately normalized by the treatment and control fractions, respectively, which corresponds to the Horvitz-Thompson estimator, or equivalently in our formal mean-field setting, the difference-in-means estimator. More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, for which the treatment approaches control: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$. We define a local version of the naive estimator as:
$\widehat{GTE}(x_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (2)
We make the following assumption that is kept throughout the paper:
{assumption}
$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $10<q<$ 1.
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that experimental estimates in practice typically implicitly rely on the assumption of constant returns to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
Now, we obtain:
$\widehat{GTE}(x_{0},q)$ $=$ $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (3)
$=$ $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{0}(x_{0},x_{0},q)/(1-q)-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
$=$ $T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q),$
The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.2. The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
Finally, we define the estimation bias of this local estimator as:
$Bias(x_{0},q)$ ::=:= $GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
$=$ $T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)$ (5)
Studying bias in this local regime proves fruitful, as it leads to the following simple characterization.
Proposition 3.1
The bias is given by the following expression:
$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q$
Proof 3.2
Proof. By Assumption 3.2, we have that:
$T(x_{0})=T_{1}(x_{0},x_{0},q)/q\ .$
By Assumption 3.2 and application of the chain rule we get:
$T^{\prime}(x_{0})=T_{1x}(x_{0},x_{0},q)/q+T_{1y}(x_{0},x_{0},q)/q\ .$
Replacing in (5), we obtain the desired result:
$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q\ .$
The result provides an intuitive characterization of bias. Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable on the other group changes marginally. The cross-partial derivatives represent the effect on the performance metric across experimental groups, making them an intuitive indicator of interference. As interference increases, bias correspondingly intensifies.
4 Revenue Management Setting
We examine a traditional revenue management scenario involving monopoly pricing. A firm or platform offers products that, at least based on observable characteristics, are deemed homogeneous and are thus priced uniformly. The firm faces an unknown to them demand function given by $D(p)$ as a function of the price $p$. The firm has a constant marginal production cost, $c\geq 0$. We make the following assumption that is kept throughout the paper and will imply Assumption 3.1 for the performance metrics we study.
{assumption}
The function $D(p)$ is differentiable and $D^{\prime}(p)<0$, for all $p\in\mathbb{R}_{+}$. At this level of abstraction, the demand function can describe a scenario with either an infinite supply or a limited supply, where a rationing mechanism determines the consumption. In a large limit model, even with supply constraints we expect the demand function to be differentiable.
To provide a concrete example with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price. That is, $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$ almost surely, where $\{x_{i}(p):i=1,...,n\}$ are i.i.d. random variables equal to one if customer $i$ buys the product at a given price $p$, and zero otherwise.
4.1 Estimands
We will consider two performance metrics of typical interest: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero). The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand) and subsequently affect profit. To achieve this goal, the firm is interested in estimating the difference in demand or profits generated when products are priced at a new “treatment” price relative to the demand or profits generated when products are priced at the status-quo “control” price. These differences correspond to the global treatment effects of interest that we derive below. To do so, we first define the “global” profit function as:
$\pi(p):=(p-c)D(p)\ .$
Following equation (3.1), we present our estimands. These are the local versions of the global treatment effects associated to demand and profits:
$GTE_{D}(p)$ ::=:= $D^{\prime}(p)$ (6)
$GTE_{\pi}(p)$ ::=:= $\pi^{\prime}(p)$ (7)
$=$ $D(p)+(p-c)D^{\prime}(p)$
Note that $GTE_{\pi}(p)$ is the marginal profit of the firm at price $p$.
4.2 Experiment: Design, Model, and Estimator
Given that the demand function is unknown, we assume the firm uses price experimentation to estimate the $GTE$s of interest. As mentioned above, canonical A/B-tests can randomize at either side of the market, and our model does not preclude either possibility. That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction $1-q$ to control. We denote by $p_{0}$ the baseline (control) price and by $p_{1}$ the treatment price. We assume that the experimental demands of control and treated subjects are given by the functions $Q_{0}(p_{0},p_{1},q)$ and $Q_{1}(p_{0},p_{1},q)$, respectively. We further define the functions:
$D_{0}(p_{0},p_{1},q)=\frac{Q_{0}(p_{0},p_{1},q)}{1-q};\quad D_{1}(p_{0},p_{1},q)=\frac{Q_{1}(p_{0},p_{1},q)}{q},$
which correspond to the demand for control and treated subjects scaled by the control and treatment fractions, respectively. Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions. Again, at this level of abstraction these demand functions could, for example, incorporate supply constraints and rationing mechanisms.
Recall that in a customer-side experiment treatment and control customers will observe all products priced at the treatment and control prices, respectively, with the proportion of customers assigned to the treatment group being $q$. In contrast, in a product-side experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1. In addition, $p_{0},p_{1}\geq c$. Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
We introduce the following additional assumption on experimental demand functions that implies Assumption 3.2. {assumption} The following holds:
  1. 1.
    $D_{0}(p_{0},p_{1},q)$ is non-decreasing in $p_{1}$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
  2. 2.
    $D_{1}(p_{0},p_{1},q)$ is non-decreasing in $p_{0}$, for all $p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
  3. 3.
    $D_{0}(p,p,q)=D_{1}(p,p,q)=D(p)$, for all $p\in\mathbb{R}_{+}$ and $10<q<$ 1.
The first two assumptions are natural as they imply that control and treatment products are (imperfect) substitutes. The third assumption states that if the control and treatment prices are the same, then both the scaled control and treatment demand functions are equivalent to the market demand function; this assumption is implied by constant returns to scale. This assumption is satisfied by demand systems for which the control and treatment demand quantities are proportional to their shares in the marketplace. Indeed, this assumption will be satisfied for the model we later study in Section 6, as well for a number of other common demand systems.
We finish this subsection by defining the experimental estimators. To do so, we define the experimental control and treatment profit functions, respectively:
$\pi_{0}(p_{0},p_{1},q)$ ::=:= $(p_{0}-c)D_{0}(p_{0},p_{1},q)$
$\pi_{1}(p_{0},p_{1},q)$ ::=:= $(p_{1}-c)D_{1}(p_{0},p_{1},q).$
Following definition (2) we define naive estimators for demand and profits associated to the canonical designs:
$\widehat{GTE}_{D}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
$\widehat{GTE}_{\pi}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
Note that analogs of these naive estimators are frequently used in practice in estimating global treatment effects from experimental data.
Now, by equation (3), we obtain:
$\widehat{GTE}_{D}(p_{0},q)$ $=$ $D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\ ,$ (8)
$\widehat{GTE}_{\pi}(p_{0},q)$ $=$ $\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$ (9)
$=$ $D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$
where we used Assumption 4.2 and the last equation follows by Assumption 4.2.3.
5 Biases: Analysis and Results
In this section we analyze biases associated to the naive estimators. Following definition (4) we define the estimation bias for demand and profits associated to the canonical designs:
$Bias_{D}(p_{0},q)$ ::=:= $GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)$
$Bias_{\pi}(p_{0},q)$ ::=:= $GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)$
First, note that by Proposition 3.1 we obtain the following bias characterizations:
$Bias_{D}(p_{0},q)$ $=$ $D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)$ (10)
$Bias_{\pi}(p_{0},q)$ $=$ $\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$ (11)
$=$ $(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$
$=$ $(p_{0}-c)Bias_{D}(p_{0},q)$
As discussed in Proposition 3.1, bias increases with increasing cross-effects between treatment and control groups. More specifically, the bias is influenced by the degree to which the control (or treatment) group detracts demand from its counterpart in response to a marginal price change in the latter group.
Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, so we obtain that the naive estimators always (weakly) underestimate $GTE$:
Proposition 5.1
We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
The result follows similar cannibalization arguments developed in previous work (Johari et al., 2022). To provide intuition consider a listing-side experiment in which treatment price is larger than control price. In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from the treatment to the control group. Similarly, experimental control demand overestimates global control demand. Now, let’s consider a customer-side experiment where there are supply constraints, so customers compete for limited products. In such cases, the control customers may “cannibalize” supply from the treatment group, given the lower control price, leading again to an underestimation of the global treatment effect. Note that in a customer-side experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
It is important to observe that because $D^{\prime}(p)<0$ by Assumption 4, $GTE_{D}(p_{0})<0$, for all $p$. Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the naive estimator will have the same (nonpositive) sign.
However, there is a more intriguing result for profits. Suppose that the price $p_{1}$ is greater than $p_{0}$. In this scenario, even if the demand $D(p_{1})$ at $p_{1}$ is lower than the demand $D(p_{0})$ at $p_{0}$, the increase in price may offset the decrease in demand, leading to a positive true global treatment effect in profits. However, because the naive estimator underestimates the treatment effect, it could be negative. Next, we indeed provide conditions under which the naive estimator not only underestimates the global treatment effect, but also obtains the wrong sign.
By definition, $GTE_{\pi}(p_{0})\geq 0$ if and only if $\pi^{\prime}(p_{0})\geq 0$, which after some simple manipulation yields the following standard “markup” inequality:
[…]
where $e_{p}(p_{0})=D^{\prime}(p_{0})p_{0}/D(p_{0})$ is the price elasticity of demand. If $\pi(p)$ is strictly quasi-concave, the condition is equivalent to the price $p_{0}$ being below the optimal monopoly price. For the case $c=0$, condition (a) naturally states that marginal revenue is increasing at $p_{0}$, or equivalently that the demand is relatively inelastic with respect to price.
Now, consider the naive estimator. Note that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ is equivalent to $Bias_{\pi}(p_{0},q)\geq{GTE}_{\pi}(p_{0})$. That is, provided that condition $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect. Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
[…]
Note that condition (b) is a “modified markup” inequality. We have the following result:
Proposition 5.2
  1. 1.
    Condition (a) holds if and only if $GTE_{\pi}(p_{0})\geq 0$.
  2. 2.
    Condition (b) holds if and only if $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$.
  3. 3.
    $A(p_{0})\leq B(p_{0},q)$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
Proof 5.3
Proof. Points (1) and (2) in the proposition are direct from the arguments above. To prove point (3), note that $D_{1y}(p_{0},p_{0},q)\leq D^{\prime}(p_{0})$ (by applying Assumptions 4.2.2 and 4.2.3 together with the chain rule) and $D_{0y}(p_{0},p_{0},q)\geq 0$ (by Assumption 4.2.1). Hence,
$\frac{p_{0}-c}{p_{0}}\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\leq\frac{p_{0}-c}{p_{0}}D^{\prime}(p_{0})\ ,$
and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4.
The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold, that is there are appropriate lower and upper bounds on the inverse of the price elasticity. Note that we use weak inequalities in conditions (a) and (b) for a change of sign; strictly speaking of course, a change of sign can only occur if at least one of the two inequalities is strict. [^2] If exactly one of these inequalities is strict (and both hold), the canonical estimator will still instruct firms to move in a non-optimal direction; either by telling them to lower prices when they were already at the optimal price, or by suggesting they keep their price constant when they were underpricing. We suppress this edge case in the sequel in favor of simplicity of presentation. [^3] In our two-sided market model, introduced in section 6, the scenario $\widehat{GTE}_{\pi}(p_{0},q)=GTE_{\pi}(p_{0})=0$ is impossible, meaning this edge case does not exist.
To provide intuition, note that by using equation (8) we can define the experimental price elasticity of demand as:
$\widehat{e_{p}}(p_{0},q)=\frac{p_{0}(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q))}{D(p_{0})}.$ (12)
Hence, we can re-write condition (b) as:
$-\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
which can be understood as the experimental analogue of a markup inequality. It states that at the experimental elasticity, the current price is above the optimal monopoly price (under an appropriate concavity assumption). Hence, as expected, a change of sign occurs when the naive estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices). The result has important consequences for experiment-driven price optimization: the prescription implied by the experiment (reduce price) is the opposite to the one implied by the $GTE$ (increase price). Hence, if one moves in the direction suggested by the experiment, profits decrease relative to the status-quo.
To provide further intuition, let us assume that condition (a) holds and that profits are quasiconcave, so that the firm is underpricing relative to the optimal monopoly price. Hence, we have a change of sign if and only if condition (b) holds. If $|D_{1y}(p_{0},p_{0},q)|+D_{0y}(p_{0},p_{0},q)$ is large compared to $|D^{\prime}(p_{0})|$, condition (b) will hold. Note that $|D_{1y}(p_{0},p_{0},q)|$ represents the magnitude of the decrease in treatment demand given a small increase in treatment price, which captures two effects: (i) demand lost to the outside option; and (ii) cannibalization to the control group. Note that (i) is the global demand effect we aim to estimate. On the other hand, the term $D_{0y}(p_{0},p_{0},q)$ represents the increase in control demand given a small increase in treatment price, which captures cannibalization from the treatment to the control group (equivalent to effect (ii)). A change of sign happens as the two cannibalization effects just mentioned become larger.
We conclude with a remark regarding our estimators, our findings on changes of sign, and guidance for practitioners. Note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit function, respectively, as is natural given our estimands. In practice, a platform will typically directly use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices. This can have some bearing in interpreting our results, depending on the sign of the difference between the treatment and control price.
Concretely, suppose that treatment involves raising the price relative to control. Then our results show that a change of sign can only occur if the true global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment. In other words, in this case, the platform risks failing to move prices in a profitable direction.
On the other hand, suppose the treatment involves lowering the price relative to control. Then our results show that a change of sign can only occur if the true global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment. In other words, in this case, the platform risks moving prices in a potentially unprofitable direction.
6 A Two-Sided Market Model
In this section, we illustrate the application of our results of the general revenue management setting introduced in Section 4 to a two-sided market model developed in Johari et al., (2022). That paper develops a stochastic dynamic market model with supply constraints in which customers book listings for a finite period of time. The paper also introduces a mean-field limit that greatly simplifies the analysis. We focus on the mean-field model here and describe its main elements.
6.1 Preliminaries
We begin by recapping the mean-field model formulation in Johari et al., (2022), adapted to our setting with price experiments. There is a two-sided platform evolving in continuous time with listings on the supply-side and customers on the demand side. There is a continuum of listings of mass $\rho$ that are homogenous with respect to their observable characteristics. Having a limited supply will play a key role in the analysis. We assume all listings have the same price $p$. At each time $t$, listings can be available or booked. The state of the system at time $t$ is given by $s_{t}$, which denotes the mass of available listings available at time $t$.
[…]
$q(s)=\frac{sv(p)}{\varepsilon+sv(p)}.$
Finally, a listing remains occupied for an expected time $\frac{1}{\tau}$ when it gets booked; in our analysis we normalize $\tau$ to be 1 without loss of generality. Note that in our model, the parameter $\lambda$ mediates market balance: when $\lambda\rightarrow 0$ the market becomes demand constrained, when $\lambda\rightarrow\infty$ the market becomes supply constrained.
Note that the general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers. In the model we consider in this section, we make both consumers and listings homogeneous.
As in Johari et al., (2022) we focus on the steady-state of the system. Let $s^{*}(\lambda,p)$ be the steady-state mass of available listings for a given market balance parameter $\lambda$ and price $p$. It is shown in Johari et al., (2022) that $s^{*}(\lambda,p)$ is the unique solution to the following conservation of flow equation:
$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)v(p)}{\varepsilon+s^{*}(\lambda,p)v(p)}:=D(\lambda,p)\ .$ (13)
The left-hand side of the equation corresponds to the rate at which listings become available. The right-hand side corresponds to the booking (or demand) rate, defined as $D(\lambda,p)$. In the steady-state both rates are equal.
{assumption}
We assume the following for the utility function $v(p)$:
  1. 1.
    $v(p)$ is strictly positive for $p\in[c,+\infty);$
  2. 2.
    $v(p)$ is strictly decreasing for $p\in[c,+\infty);$
  3. 3.
    $v(p)$ is differentiable for $p\in[c,+\infty)$, and $v^{\prime}(p)$ is continuous;
  4. 4.
    $-pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$
[…]
6.2 Steady-State and Estimand
To gain some initial intuition, we start this section by providing some analysis on the steady-state and the associated demand. Then, we define our estimands.
First, we note that listing availability $s^{*}(\lambda,p)$ is monotonically increasing with price $p$ and decreasing with the market balance parameter $\lambda$, which is natural. These monotonicity results can be derived by differentiating $s^{*}(\lambda,p)$ in equation (13). [^7] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\beta,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$. Figure 4 in the Appendix illustrates the steady-state listing availability as a function of price ($p$) and market balance ($\lambda$).
Naturally, $D(\lambda,p)$ is monotonically decreasing in price ($p$) and increasing in market balance ($\lambda$). Again, we can prove these monotonicity results by differentiating. [^8] This can be seen by rewriting the demand as the rate at which the listings become available. Using the market balance equation, we have $D(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)v(p)}{\varepsilon+s^{*}(\lambda,p)v(p)}=\rho-s^{*}(\lambda,p)$. Hence, $D_{p}(\lambda,p)=-s^{*}_{p}(\lambda,p)$ and $D_{\lambda}(\lambda,p)=-s^{*}_{\lambda}(\lambda,p)$. Since $s^{*}_{p}(\lambda,p)>0$ and $s^{*}_{\lambda}(\lambda,p)<0$ we obtain that $D_{p}(\lambda,p)<0$ and $D_{\lambda}(\lambda,p)>0$. Figure 5 in the Appendix shows the dependence of the demand function on the market balance parameter and price.
Now, we define our estimands. Following the definitions in equations (6) and (7):
$GTE_{D}(\lambda,p)$ $=D_{p}(\lambda,p)$
$GTE_{\pi}(\lambda,p)$ $=D(\lambda,p)+(p-c)D_{p}(\lambda,p)$
$=\rho-s^{*}(\lambda,p)-(p-c)s^{*}_{p}(\lambda,p),$ (14)
where in the last equation we used equation (13).
6.3 Experimental Designs
In this section we analyze two canonical experimental designs in platforms: randomizing listings or randomizing customers.
6.3.1 Listing-Side Randomization
We assume that randomization occurs on the listing side, where a fraction $q$ of the listings is assigned to treatment. Hence, a customer is exposed to a mix of treatment and control listings. In this experimental design, the balance equations are given by:
$(1-q)\rho-s_{0}^{*}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})+s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}$ (15)
$q\rho-s_{1}^{*}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s_{0}^{*}(\lambda,p_{0},p_{1},q)v(p_{0})+s_{1}^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}\ ,$
where $s^{*}_{0}(\lambda,p_{0},p_{1},q)$ and $s^{*}_{1}(\lambda,p_{0},p_{1},q)$ refer to the steady-state available control and treated listings, respectively.
The control and treatment demand functions $D^{LR}_{0}(\lambda,p_{0},p_{1},q)$ and $D_{1}^{LR}(\lambda,p_{0},p_{1},q)$ are given by:
$D_{0}^{LR}(\lambda,p_{0},p_{1},q)$ $=\rho-s_{0}^{*}(\lambda,p_{0},p_{1},q)/(1-q);$ (16)
$D^{LR}_{1}(\lambda,p_{0},p_{1},q)$ $=\rho-s_{1}^{*}(\lambda,p_{0},p_{1},q)/q.$
We can show that $D^{LR}_{0}(\lambda,p_{0},p_{1},q)$ and $D^{LR}_{1}(\lambda,p_{0},p_{1},q)$ satisfy Assumption 4.2 (see Appendix 10.4).
Following the characterization of bias from equations (10) and (11), we obtain that the bias in this setting, $Bias_{\pi}^{LR}(\lambda,p_{0},q)$ , is given by:
$Bias^{LR}_{\pi}(\lambda,p_{0},q)$ $=(p_{0}-c)(D^{LR}_{1x}(\lambda,p_{0},p_{0},q)+D^{LR}_{0y}(\lambda,p_{0},p_{0},q))$
$=(p_{0}-c)\frac{-\lambda s^{*}(\lambda,p_{0})^{2}v(p_{0})v^{\prime}(p_{0})(\varepsilon+s^{*}(\lambda,p_{0})v(p_{0}))}{[\varepsilon+s^{*}(\lambda,p_{0})v(p_{0})+\lambda v(p_{0})][(s^{*}(\lambda,p_{0})v(p_{0})+\varepsilon)^{2}+\lambda v(p_{0})\varepsilon]}$ (17)
Note that the bias is independent of the treatment fraction $q$.
6.3.2 Customer-Side Randomization
In this case we randomize on the customer side, with a fraction $q$ of customers randomized into treatment. In this case, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. The steady-state is given by:
$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\lambda\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (18)
In this setting we have the following demand functions:
$D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})},$ (19)
$D^{CR}_{1}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}.$
One can show that the demand functions $D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ and $D^{CR}_{1}(\lambda,p_{0},p_{1},q)$ satisfy Assumption 4.2 (see Appendix 10.5).
Using the characterization in equation (11) we obtain the following expression for $Bias_{\pi}^{CR}$:
$Bias^{CR}_{\pi}(\lambda,p_{0},q)$ $=(p_{0}-c)(D^{CR}_{1x}(\lambda,p_{0},p_{0},q)+D^{CR}_{0y}(\lambda,p_{0},p_{0},q))$
$=-(p_{0}-c)\frac{\lambda^{2}\varepsilon^{2}v(p_{0})v^{\prime}(p_{0})s^{*}(\lambda,p_{0})}{(v(p_{0})s^{*}(\lambda,p_{0})+\varepsilon)^{2}\left[(s^{*}(\lambda,p_{0})v(p_{0})+\varepsilon)^{2}+\lambda v(p_{0})\varepsilon\right]}$ (20)
Note that the bias in the CR-setting is independent of the treatment fraction $q$.
6.4 Properties of Biases with Market Balance
In this section we study structural properties of biases with a particular focus on extreme regimes of market balance. The analysis in this subsection closely follows Johari et al., (2022).
[…]
As the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines resulting in less interference in a CR experiment. In the limit, there is no competition between customers and we expect the bias in the CR experiment to be zero. On the other hand, competition between listings increases resulting in more interference in a LR experiment. In the limit, we expect the bias in the LR experiment to be non-zero.
Conversely as the market becomes supply constrained ($\lambda\rightarrow\infty$), competition between listings declines resulting in less interference in a LR experiment. In the limit, there is no competition between listings and we expect the bias in the LR experiment to be zero. Competition between customers increases resulting in more interference in a CR experiment. In the limit, we expect the bias in the CR experiment to be non-zero.
The following propositions formalize the intuition about the behavior of the bias in both extremes of market balance.
Proposition 6.1
For all $p\in[c,+\infty)$, we have that:
$\lim_{\lambda\to 0}GTE_{D}(\lambda,p)/\lambda=\frac{\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}};\ \ \lim_{\lambda\to 0}GTE_{\pi}(\lambda,p)/\lambda=\frac{\rho v(p)}{\varepsilon+\rho v(p)}+\frac{(p-c)\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}}.$
We have that $\lim_{\lambda\to 0}Bias_{\pi}^{CR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{CR}(\lambda,p)/\lambda=0$, while
$\lim_{\lambda\to 0}Bias_{\pi}^{LR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{LR}(\lambda,p)/\lambda=-\frac{(p-c)\rho^{2}v(p)v^{\prime}(p)}{(\varepsilon+\rho v(p))^{2}}.$
Proposition 6.2
For all $p\in[c,+\infty)$, we have that: $\lim_{\lambda\to\infty}GTE_{D}(\lambda,p)=0$ and that:
$\lim_{\lambda\to\infty}GTE_{\pi}(\lambda,p)=\rho$. We have $\lim_{\lambda\to\infty}Bias_{\pi}^{LR}(\lambda,q)=(p-c)\lim_{\lambda\to\infty}Bias_{D}^{LR}(\lambda,q)=0$, while
$\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=(p-c)\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=-\frac{\rho(p-c)v^{\prime}(p)}{v(p)}.$
In the limit where $\lambda\rightarrow 0$, Proposition 6.1 suggests that the CR experiment is unbiased while the LR experiment is biased. Similarly, in the limit where $\lambda\rightarrow\infty$, Proposition 6.2 suggests that the CR experiment is biased while the LR experiment is unbiased. Proofs of both propositions are presented in Appendix 10. The proofs are similar to the analysis in Johari et al., (2022) but we also leverage our characterization of biases based on cross-partial derivatives. In the next section we will further study the extent of these biases to characterize change of sign regimes.
We also provide some figures to complement our limit results. In Figure 1. As $\lambda$ increases, the market becomes more supply-constrained and bias decreases as competition between listings dissipates. Figure 2 shows the normalized bias of profits for CR, and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$. As the price $p$ changes, the sensitivity of demand to price changes as well; in particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit.
Figure 1: Illustration of $Bias^{LR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$). Instance is given by $v(p)=e^{V-p}$, $\epsilon=1$, $c=1$, $V=5$, and $\rho=1$.
Figure 2: Illustration of $Bias^{CR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$). Instance is given by $v(p)=e^{V-p}$, $\epsilon=1$, $c=1$, $V=5$, and $\rho=1$.
6.5 Change of Sign
We recall the two conditions for a change of sign to occur:
[…]
An interesting immediate observation is that the regimes in which a change in sign occurs in an LR or CR experiment using the corresponding naive estimators are independent of the treatment fraction $q$, because the corresponding biases are independent of $q$. This has important implications for practice, as the treatment fraction is not a lever that can prevent the possibility that the sign of the treatment effect estimate is incorrect.
Using the structure of the model we can provide the following insightful characterizations for the change of sign regions, as follows.
Theorem 6.3
For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the listing-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{LR}_{\lambda}$.
Theorem 6.4
For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the customer-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{CR}_{\lambda}$.
Theorems 6.3 and 6.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run. The proofs are provided in Appendix 10.9 and repeatedly exploit the market balance equations for CR and LR respectively.
The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$. In particular, this reveals that the sets $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not comprised of unrealistic prices: if a firm is currently pricing optimally or very slightly below optimally, the sign of their canonical estimator will be wrong, regardless of their experimental setup.
It is not true, however, that the two experimental setups are equally fallible in all settings. Note that $\widehat{e_{p}}^{LR}$ is not equal to $\widehat{e_{p}}^{CR}$, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same, and behave quite differently based on market balance, as we illustrate through the following numerical study.
6.5.1 Numerical Results
Figure (3) shows numerical results on the set of prices where a change of sign occurs.
Figure 3: Illustration of the change of sign region (white) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in black. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in grey.
Consistent with Propositions 6.1 and 6.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment. At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
7 An Empirical Case Study
Our results show that price experiments can get the sign wrong, and suggest that practitioners would benefit from investigating whether their experiments are vulnerable to such changes of sign. In this section, we present an empirical illustration to study this question, by calibrating a version of our stylized model to data from Airbnb. Although our model simplifies the complexity of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with empirical data. As we will see, our results suggest that a change of sign is indeed plausible in the Airbnb context.
We rely on the estimates by Holtz et al., (2020) of a guest fee “meta-experiment” at Airbnb. This experiment involved a treatment that raised platform fees. The design consisted of two “arms”: randomization across clusters of listings in one arm; as well as Bernoulli randomization across listings in a second arm. Holtz et al., (2020) use this meta-experiment to estimate the degree to which Bernoulli randomized experiments exhibit interference, by comparing with estimates from the cluster randomized arm (which mitigates interference effects). The paper reports on cluster and naive experiment estimates and standard errors for the global treatment effect (the TATE, or total average treatment effect in their paper) for bookings, nights booked, and gross spend.
To translate from Holtz et al., (2020) to our setting, we first recap certain features of their presentation. First, even though the treatment corresponds to a percentage increase in prices, the size of this treatment intervention is not disclosed in Holtz et al., (2020). For our illustrative exercise here, we interpret the estimates of Holtz et al., (2020) as semi-elasticities, i.e., absolute changes in quantities for a 1% price increase. Second, Holtz et al., (2020) does not disclose absolute levels (e.g., of nights booked, or revenue); instead, these quantities are rescaled by the same randomly chosen quantity (not disclosed to the reader). As we will see, this does not present any challenges for our exercise here.
We need to slightly adjust the analysis to accommodate a percentage change in prices relative to the bias analysis in Section 5 that considers an absolute change. Such an analysis can be carried out by considering demand $D(\alpha p)$ and profit $(\alpha p-c)D(\alpha p)$, and considering derivatives with respect to $\alpha$, evaluated at $\alpha=1$. Proceeding in this way, we redefine:
$GTE_{D}(p)=pD^{\prime}(p);\ \ GTE_{\pi}(p)=p(D(p)+pD^{\prime}(p))-cpD^{\prime}(p)$
as the semi-elasticities of demand and profit, respectively.
As before, the first condition for a change of sign at $p_{0}$ is (a) $GTE_{\pi}(p_{0})\geq 0$, which becomes:
$p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))-cp_{0}D^{\prime}(p_{0}).$ (21)
The second condition is that (b) $Bias_{\pi}(p_{0},q)\geq GTE_{\pi}(p_{0})$ (where now we note that the bias is computed using percentage price changes, i.e., with respect to changes in $\alpha$). Since $Bias_{\pi}(p_{0},q)=(p_{0}-c)Bias_{D}(p_{0},q)$, condition (b) becomes equivalent to:
$(p_{0}-c)Bias_{D}(p_{0},q)\geq p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))p_{0}-cp_{0}D^{\prime}(p_{0}).$ (22)
We now leverage the empirical estimates of Holtz et al., (2020). The cluster arm estimates are a decrease of $1.136$ in nights booked per listing, and of $116.67$ in gross revenues. For our analysis, we assume that the cluster arm provides ground truth, and the corresponding bias for demand is reported to be $0.205$. We assume that $D(p_{0})$ represents nights booked per listing and $p_{0}$ the average price per night, so that $p_{0}D(p_{0})$ represents gross revenues per listing. In particular, note that then $p_{0}D^{\prime}(p_{0})$ is the semi-elasticity of nights booked, and $p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))$ is the semi-elasticity of gross revenues per listing. Finally, we assume $c$ represents the marginal (opportunity) cost per night.
By equation (21) and the estimates, the first condition for change of sign amounts to $-116.7+1.136c\geq 0$, because we identify $-116.67$ and $-1.136$ as the ground-truth semi-elasticities of revenues and demand, respectively. This simplifies to
$c\geq 102.73.$ (23)
The second condition for change of sign amounts to $0.205(p_{0}-c)\geq-116.67+1.136c$, which can be simplified to
$116.67+0.205p_{0}\geq 1.341c$ (24)
Importantly, note that all quantities used in these inequalities are re-scaled by the same random constant, and so these inequalities are valid for any value of that constant under our assumptions.
The pair of inequalities (23)-(24) can be used to interrogate whether a change of sign is plausible in this setting. For example, for a reasonable marginal opportunity cost of $c=125$ (corresponding to \$125 per night), our analysis suggests that market segments with a baseline price of $p_{0}=249$ (corresponding to \$249 per night) yield an incorrectly estimated sign. More generally, this exercise allows us to identify combinations of prices and costs for which the canonical estimator would result in a change of sign. In particular, (24) can be rewritten as $0.847-87.002/p_{0}\leq(p_{0}-c)/p_{0}$, lower bounding markups based on observed market prices that yield a change of sign. For example, the average price of recently booked Airbnb listings in San Francisco is given by \$212 (cf http://insideairbnb.com/san-francisco). According to our analysis, this implies a markup of $0.43$ or larger would yield a change of sign. Given that these are markups that are reasonable in practice, we believe this stylized exercise motivates further detailed empirical study of the prevalence of such changes of sign in real-world platforms.
8 Conclusion and Future Work
In this paper we studied biases that arise due to interference in platform experiments, with a particular emphasis on profit or revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue.
In future work, we aim to extend the analysis to other settings of interest, including two-sided market settings where the variable of interest is the percentage fee charged by the platform (as a fraction of gross revenue). Further, our findings in Section 7 motivate further empirical study of the prevalence of changes of sign in practice. It is natural to consider more robust empirical calibration strategies, as well as practical methodological approaches to mitigating this change of sign. These remain important directions for future work.

### v3

@@
 Price Experimentation and Interference
+[…]
 […]
@@
 In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo.
-We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup.Finally, we discuss structural and practical implications for platform operators.
+We introduce a novel debiasing technique for such experiments, that only requires that the firm equally split experimental units between treatment and control.
+We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup.
+We also extend our revenue management model to an alternative specification in which firms set a fixed sales fee across heterogeneous products, and show that all our results extend to this setting as well.
+[…]
 […]
@@
 Our results have important implications for price optimization based on experiments: incorrectly following the experiment’s suggestion may lead one to move in the wrong direction and decrease profits relative to the status-quo!
-We apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022).
+We take advantage of our structural characterization of bias to construct a novel debiasing method for the canonical A/B experiment.
+In particular, we show that if the experiment equally allocates units between treatment and control, then it is possible to fully debias the canonical estimator, by taking advantage of symmetry in the structure of bias.
+Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below).
+This is a stark contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require much more complicated experimental designs and/or estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
+To obtain additional insight, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022).
 We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs.
-For example, for a listing-side experiment a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low.
-We also calibrate this stylized model to experimental estimates from Airbnb (published in Holtz et al., 2020), and investigate empirical implications for regimes where a change of sign may occur in a price experiment.
+For example, for a listing-side experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low.
+We also investigate numerically how bias depends on both market balance and the baseline control price.
+We conclude the paper with an alternative specialization of our revenue management setting to a model where the firm sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces.
+In this model we allow the possibility of a heterogeneous collection of goods for sale, with an unknown (joint) demand function.
+We show that all our previous results extend to this setting as well.
 2 Related Work
@@
 Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice.
-Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, farias2023correcting), and message passing (shirani2023causal).
+Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), and message passing (shirani2023causal).
 A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023).
 Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects).
-roemheld2023interference uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments.
+Roemheld and Rao, (2024) uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments.
 More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods.
@@
 $=$ $(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$
-$=$ $(p_{0}-c)Bias_{D}(p_{0},q)$
+$=$ $(p_{0}-c)Bias_{D}(p_{0},q)$ (12)
 As discussed in Proposition 3.1, bias increases with increasing cross-effects between treatment and control groups.
@@
 In other words, in this case, the platform risks moving prices in a potentially unprofitable direction.
+6 Debiasing Canonical Estimators
+In this section, we take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that applies as long as the platform has equally allocated units between treatment and control, i.e., $q=1/2$.
+Recall (10) (11) provide the expressions for bias in terms of derivatives of demand.
+We now show how we can use observables to estimate this bias.
+First, we show that $D_{0y}(p_{0},p_{0},q)$ can be computed (approximately) using observed data.
+To see this, note that $D_{0y}(p_{0},p_{0},q)$ is defined by:
+$D_{0y}(p_{0},p_{0},q)=\lim_{p_{1}\to p_{0}^{+}}\frac{D_{0}(p_{0},p_{1},q)-D_{0}(p_{0},p_{0},q)}{p_{1}-p_{0}}.$
+By Assumption 6 we have that $D_{0}(p_{0},p_{0},q)=D(p_{0})$.
+Since we assume that the firm was pricing at $p_{0}$ before the experiment began, they have observed $D(p_{0})$ prior to the experiment.
+The quantity $D_{0}(p_{0},p_{1},q)$ is observed as well: it represents the control demand during the experiment.
+This means that in the limit as $p_{1}$ approaches $p_{0}$, observables can be used to approximately compute the derivative $D_{0y}(p_{0},p_{0},q)$.
+We now note that by symmetry, we have $D_{0}(p_{0},p_{1},1/2)=D_{1}(p_{1},p_{0},1/2)$.
+This holds because both terms represent demand for the half of the goods that are priced at $p_{0}$, when the other half of goods are priced at $p_{1}$.
+Taking the derivative of both sides of this equality with respect to $p_{1}$, we obtain $D_{0y}(p_{0},p_{1},1/2)=D_{1x}(p_{1},p_{0},1/2)$, so that when $p_{1}=p_{0}$, there holds $D_{0y}(p_{0},p_{0},1/2)=D_{1x}(p_{0},p_{0},1/2)$.
+We have shown the following result.
+Proposition 6.1
+Suppose that $q=1/2$. Then there holds:
+$Bias_{D}(p_{0},1/2)$ $=2D_{0y}(p_{0},p_{0},1/2);$ (14)
+$Bias_{\pi}(p_{0},1/2)$ $=2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (15)
+In particular, there holds:
+$GTE_{D}$ $=\widehat{GTE}_{D}(p_{0},1/2)+2D_{0y}(p_{0},p_{0},1/2);$ (16)
+$GTE_{\pi}$ $=\widehat{GTE}_{\pi}(p_{0},1/2)+2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (17)
+Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables, then the right hand sides of (16) and (17) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
+We note several limitations of this method.
+First, of course, the symmetry property being exploited only applies when the market is equally allocated to treatment and control.
+Second, the assumption that $D(p_{0})$ is observed relies on market stationarity; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment.
+Finally, estimation of $D_{0y}$ uses only half of the experimental data, namely, only the observed control demand, which may be a source of sample inefficiency (i.e., increased variance).
 7 A Two-Sided Market Model
@@
 These monotonicity results can be derived by differentiating $s^{*}(\lambda,p)$ in equation (18).
-[^7] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\beta,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$.
+[^7] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\lambda,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$.
 Figure 4 in the Appendix illustrates the steady-state listing availability as a function of price ($p$) and market balance ($\lambda$).
@@
 The steady-state is given by:
-$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\lambda\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (18)
+$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (23)
 In this setting we have the following demand functions:
@@
 For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the customer-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{CR}_{\lambda}$.
-Theorems 6.3 and 6.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run.
+Theorems 7.3 and 7.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run.
 The proofs are provided in Appendix 11.9 and repeatedly exploit the market balance equations for CR and LR respectively.
@@
 The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in grey.
-Consistent with Propositions 6.1 and 6.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment.
+Consistent with Propositions 7.1 and 7.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment.
 At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest.
@@
 We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
-7 An Empirical Case Study
-Our results show that price experiments can get the sign wrong, and suggest that practitioners would benefit from investigating whether their experiments are vulnerable to such changes of sign.
-In this section, we present an empirical illustration to study this question, by calibrating a version of our stylized model to data from Airbnb.
-Although our model simplifies the complexity of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with empirical data.
-As we will see, our results suggest that a change of sign is indeed plausible in the Airbnb context.
-We rely on the estimates by Holtz et al., (2020) of a guest fee “meta-experiment” at Airbnb.
-This experiment involved a treatment that raised platform fees.
-The design consisted of two “arms”: randomization across clusters of listings in one arm; as well as Bernoulli randomization across listings in a second arm.
-Holtz et al., (2020) use this meta-experiment to estimate the degree to which Bernoulli randomized experiments exhibit interference, by comparing with estimates from the cluster randomized arm (which mitigates interference effects).
-The paper reports on cluster and naive experiment estimates and standard errors for the global treatment effect (the TATE, or total average treatment effect in their paper) for bookings, nights booked, and gross spend.
-To translate from Holtz et al., (2020) to our setting, we first recap certain features of their presentation.
-First, even though the treatment corresponds to a percentage increase in prices, the size of this treatment intervention is not disclosed in Holtz et al., (2020).
-For our illustrative exercise here, we interpret the estimates of Holtz et al., (2020) as semi-elasticities, i.e., absolute changes in quantities for a 1% price increase.
-Second, Holtz et al., (2020) does not disclose absolute levels (e.g., of nights booked, or revenue); instead, these quantities are rescaled by the same randomly chosen quantity (not disclosed to the reader).
-As we will see, this does not present any challenges for our exercise here.
-We need to slightly adjust the analysis to accommodate a percentage change in prices relative to the bias analysis in Section 5 that considers an absolute change.
-Such an analysis can be carried out by considering demand $D(\alpha p)$ and profit $(\alpha p-c)D(\alpha p)$, and considering derivatives with respect to $\alpha$, evaluated at $\alpha=1$.
-Proceeding in this way, we redefine:
-$GTE_{D}(p)=pD^{\prime}(p);\ \ GTE_{\pi}(p)=p(D(p)+pD^{\prime}(p))-cpD^{\prime}(p)$
-as the semi-elasticities of demand and profit, respectively.
-As before, the first condition for a change of sign at $p_{0}$ is (a) $GTE_{\pi}(p_{0})\geq 0$, which becomes:
-$p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))-cp_{0}D^{\prime}(p_{0}).$ (21)
-The second condition is that (b) $Bias_{\pi}(p_{0},q)\geq GTE_{\pi}(p_{0})$ (where now we note that the bias is computed using percentage price changes, i.e., with respect to changes in $\alpha$).
-Since $Bias_{\pi}(p_{0},q)=(p_{0}-c)Bias_{D}(p_{0},q)$, condition (b) becomes equivalent to:
-$(p_{0}-c)Bias_{D}(p_{0},q)\geq p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))p_{0}-cp_{0}D^{\prime}(p_{0}).$ (22)
-We now leverage the empirical estimates of Holtz et al., (2020).
-The cluster arm estimates are a decrease of $1.136$ in nights booked per listing, and of $116.67$ in gross revenues.
-For our analysis, we assume that the cluster arm provides ground truth, and the corresponding bias for demand is reported to be $0.205$.
-We assume that $D(p_{0})$ represents nights booked per listing and $p_{0}$ the average price per night, so that $p_{0}D(p_{0})$ represents gross revenues per listing.
-In particular, note that then $p_{0}D^{\prime}(p_{0})$ is the semi-elasticity of nights booked, and $p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))$ is the semi-elasticity of gross revenues per listing.
-Finally, we assume $c$ represents the marginal (opportunity) cost per night.
-By equation (21) and the estimates, the first condition for change of sign amounts to $-116.7+1.136c\geq 0$, because we identify $-116.67$ and $-1.136$ as the ground-truth semi-elasticities of revenues and demand, respectively.
-This simplifies to
-$c\geq 102.73.$ (23)
-The second condition for change of sign amounts to $0.205(p_{0}-c)\geq-116.67+1.136c$, which can be simplified to
-$116.67+0.205p_{0}\geq 1.341c$ (24)
-Importantly, note that all quantities used in these inequalities are re-scaled by the same random constant, and so these inequalities are valid for any value of that constant under our assumptions.
-The pair of inequalities (23)-(24) can be used to interrogate whether a change of sign is plausible in this setting.
-For example, for a reasonable marginal opportunity cost of $c=125$ (corresponding to \$125 per night), our analysis suggests that market segments with a baseline price of $p_{0}=249$ (corresponding to \$249 per night) yield an incorrectly estimated sign.
-More generally, this exercise allows us to identify combinations of prices and costs for which the canonical estimator would result in a change of sign.
-In particular, (24) can be rewritten as $0.847-87.002/p_{0}\leq(p_{0}-c)/p_{0}$, lower bounding markups based on observed market prices that yield a change of sign.
-For example, the average price of recently booked Airbnb listings in San Francisco is given by \$212 (cf http://insideairbnb.com/san-francisco).
-According to our analysis, this implies a markup of $0.43$ or larger would yield a change of sign.
-Given that these are markups that are reasonable in practice, we believe this stylized exercise motivates further detailed empirical study of the prevalence of such changes of sign in real-world platforms.
+8 Fee Model
+In this section we consider an environment in which a firm charges a fixed sales fee, as opposed to setting a price and incurring a marginal cost for each purchase.
+This represents the common situation–particularly in two-sided online marketplaces–where firms do not directly control the prices of the goods on their platforms, but instead collect a fixed fee for the sales of heterogeneous goods.
+8.1 Specifications
+Consider $N$ products, with a fixed vector of prices $\vec{p}\in\mathbb{R}^{N}$.
+The demand function $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$, takes in a vector of prices and outputs a vector in $\mathbb{R}^{N}$ that represents the quantity demanded for each good.
+We assume that the platform selects a single fee for all products $f\in\mathbb{R}^{+}$; this fee represents a percentage of all money transacted that the platform will collect.
+We assume that this fee is passed on entirely to the customer and that the platform’s marginal cost is zero.
+Therefore, we can write the platform’s profit $\Pi(f,\vec{p})$ as follows: [^10] Prices may be a function of the fee, since firms may pass-through fee changes.
+However, since experiments tend to be short-lived we ignore these longer term calibration effects.
+$\Pi(f,\vec{p})=f\vec{p}\cdot D((1+f)\vec{p}).$ (26)
+Let $D^{\prime}(\cdot)$ represent the matrix of demand derivatives in price. We can then write the platform’s profit derivative in the fee as:
+$\Pi_{f}(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p})+f\vec{p}\cdot\vec{p}\cdot D^{\prime}((1+f)\vec{p}).$ (27)
+Define $R(f,\vec{p})$ as the total money transacted in the platform when the fee is $f$ and the prices are $\vec{p}$. We can then write $R(f,\vec{p})$ as
+$R(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p}).$
+Let $R_{f}(f,\vec{p})$ represent the derivative of this quantity with respect to the fee. We can then rewrite (27) as:
+$\Pi_{f}(f,\vec{p})=R(f,\vec{p})+fR_{f}(f,\vec{p}).$ (28)
+And rewrite (26) as:
+$\Pi(f,\vec{p})=fR(f,\vec{p}).$
+{assumption}
+We assume that $\Pi_{f}$ exists and is well defined for all $f\in\mathbb{R}^{+}$ and all $\vec{p}\in\mathbb{R}^{N}$, or in other words, that $\Pi$ is differentiable over its domain.
+When running an A/B experiment, the platform will compare a control fee $f_{0}$ to a treatment fee $f_{1}$.
+To keep the presentation succinct, in this section we focus only on estimation of profits.
+This would cause a true global treatment effect (GTE) of $\Pi(f_{1},\vec{p})-\Pi(f_{0},\vec{p})$.
+Thus we define the (infinitesimal) $GTE$ at the control fee $f_{0}$ as:
+$GTE_{\Pi}(f_{0},\vec{p})=\Pi_{f}(f_{0},\vec{p}).$ (29)
+We assume that a fraction $q$ of all subjects are put in treatment and $1-q$ into control.
+Note that as in the monopoly pricing example, this randomization can occur on either the customer or product.
+For simplicity, in product side randomization, we assume that for each of the $n$ types of products, a fraction $q$ of each is in treatment treatment.
+Let $\Pi_{0}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from control subjects during the experiment, and $\Pi_{1}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from treatment subjects.
+Let $R_{0}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for control goods, and $R_{1}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for treatment goods.
+We then write $\Pi_{0}(f_{0},f_{1},\vec{p},q)=f_{0}R_{0}(f_{0},f_{1},\vec{p},q)$ and $\Pi_{1}(f_{0},f_{1},\vec{p},q)=f_{1}R_{1}(f_{0},f_{1},\vec{p},q)$.
+We make the following assumptions for these functions, analogous to Assumption 4.2.
+{assumption}
+The following holds:
+  1. 1.
+    $R_{0}(f_{0},f_{1},\vec{p},q)$ and $R_{1}(f_{0},f_{1},\vec{p},q)$ are both differentiable over their domain.
+  2. 2.
+    $R_{1}(f,f,\vec{p},q)/q=R_{0}(f,f,\vec{p},q)/(1-q)=R(f,\vec{p}).$
+Let $f_{1}^{n}$ be a sequence such that $\lim_{n\to\infty}f_{1}^{n}=f_{0}$. We then write the canonical estimator $\widehat{GTE}_{\Pi}$ as follows:
+$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\lim_{n\to\infty}\frac{\Pi_{1}(f_{0},f_{1}^{n},\vec{p},q)/q-\Pi_{0}(f_{0},f_{1}^{n},\vec{p},q)/(1-q)}{f_{1}^{n}-f_{0}}.$
+This is equivalent to:
+$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\Pi_{1y}(f_{0},f_{0},\vec{p},q)/q-\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q),$ (30)
+where $\Pi_{0y}$ and $\Pi_{1y}$ represent the derivative of the control and treatment platform profit functions with respect to the second argument respectively.
+We define the bias as $Bias_{\Pi}(f_{0},\vec{p},q)=GTE_{\Pi}(f_{0},\vec{p},q)-\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)$.
+Using Assumption 8.1 we get that $\Pi(f_{0},\vec{p},q)=\Pi_{1}(f_{0},f_{0},\vec{p},q)/q$.
+We can then use an analogous argument to our derivation of (11) to write bias as:
+$Bias_{\Pi}(f_{0},\vec{p},q)=\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+\Pi_{1x}(f_{0},f_{0},\vec{p},q)/q.$ (31)
+Using out revenue functions and equation (28) we can rewrite the preceding expressions to be more useful. First we rewrite bias as:
+$Bias_{\Pi}(f_{0},\vec{p},q)=f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+f_{0}R_{1x}(f_{0},f_{0},\vec{p},q)/q.$
+Finally, we rewrite $\widehat{GTE}_{\Pi}$ from equation (30) as follows:
+$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=R_{1}(f_{0},f_{0},\vec{p},q)/q+f_{0}R_{1y}(f_{0},f_{0},\vec{p},q)/q-f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q).$
+As before, there are two necessary and sufficient conditions for the naive estimator to have the wrong sign in a fee experiment.
+Note that as was true in the pricing problem, bias is always positive when firms test a small price increase.
+This means the general conditions for a sign flip to occur are $GTE(f_{0},\vec{p})\geq 0$, meaning the true treatment effect of a small price increase is non-negative.
+And $Bias(f_{0},\vec{p},q)\geq\widehat{GTE}(f_{0},\vec{p},q)$, meaning there is enough bias to cause the estimated treatment effect to be non-positive.
+These two conditions can be expressed as follows. First:
+[…]
+And second:
+[…]
 9 Conclusion and Future Work
@@
 Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue.
-In future work, we aim to extend the analysis to other settings of interest, including two-sided market settings where the variable of interest is the percentage fee charged by the platform (as a fraction of gross revenue).
-Further, our findings in Section 7 motivate further empirical study of the prevalence of changes of sign in practice.
-It is natural to consider more robust empirical calibration strategies, as well as practical methodological approaches to mitigating this change of sign.
-These remain important directions for future work.
+We then show a novel method for debiasing canonical A/B experiments, by splitting half of subjects into treatment and half into control.
+In future work we aim to extend our debiasing result to allow for correcting canonical experiments beyond the equal split between control and treatment.
+We also aim to run empirical studies to get a better sense of the prevalence of changes of sign in practice.

### v4

@@
 […]
-In this paper, we examine biases arising in A/B tests where firms modify a continuous parameter, such as price, to estimate the global treatment effect of a given performance metric, such as profit.
-These biases emerge in canonical experimental estimators due to interference among market participants.
+In this paper, we examine the biases that arise when firms run A/B tests on continuous parameters to estimate global treatment effects on performance metrics of interest; we particularly focus on price experiments to measure the price impact on quantity demanded, and on profit.
+In canonical A/B experimental estimators, biases emerge due to interference between market participants.
 We employ structural modeling and differential calculus to derive intuitive characterizations of these biases.
-We then specialize our general model to a standard revenue management pricing problem.
-This setting highlights a key pitfall in the use of A/B pricing experiments to guide profit maximization: notably, the canonical estimator for the expected change in profits can have the wrong sign.
-In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo.
-We introduce a novel debiasing technique for such experiments, that only requires that the firm equally split experimental units between treatment and control.
-We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup.
-We also extend our revenue management model to an alternative specification in which firms set a fixed sales fee across heterogeneous products, and show that all our results extend to this setting as well.
-[…]
-[…]
+We then specialize our general model to the standard revenue-management pricing problem.
+This setting highlights a fundamental risk innate to A/B pricing experiments: that the canonical estimator for the expected change in profits, counterintuitively, can have the wrong sign in expectation.
+In other words, following the guidance of canonical estimators may lead firms to move prices (or fees) in the wrong direction, inadvertently decreasing profits.
+We introduce a novel debiasing technique for these canonical experiments, requiring only that firms equally split units between treatment and control.
+We apply these results to a two-sided market model, and demonstrate how the “change of sign” regime depends on market factors such as the supply/demand imbalance, and the price markup.
+We conclude by calibrating our two-sided market model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners.
+[^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his critical assistance for our Airbnb empirical case study.
+We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS Revenue Management and Pricing Conferencea, the INFORMS Annual Meeting, the INFORMS Manufacturing and Service Operations Management Conference, and Airbnb for helpful feedback.
+Orrie Page was supported by a National Science Foundation Graduate Research Fellowship.
 […]
@@
 1 Introduction
-In this work we analyze the biases that arise when a technological firm or platform conducts A/B tests by varying a continuous parameter to determine the global treatment effect associated to a performance metric of interest.
-These biases appear when using canonical designs and estimators and result from interference between market participants.
-More concretely, consider the example of a canonical pricing experiment where a fraction of the products are assigned a price that is higher than the baseline price.
+In this work, we analyze the biases that arise when a firm or platform conducts an A/B test by varying a continuous parameter, in order to determine the global treatment effect associated with a performance metric of interest.
+These biases appear when using canonical experimental designs and estimators, and are the result of interference between market participants.
+More concretely, consider the example of a typical pricing experiment, commonly used by online platforms, where a fraction of the products are assigned a treatment price that is higher than the baseline (control) price.
 The global treatment effect of this higher price compares two counterfactual universes: one where all products have the high price (global treatment), and one where all products have the low price (global control).
-Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price; when using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect.
+Observe that in the experiment, some consumers who would have purchased at the high price in global treatment instead purchase control units at the low price.
+When using canonical estimators (such as the difference-in-means estimator), this substitution within the experiment biases the estimation of the global treatment effect.
 This phenomenon is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts; see Related Work (Section 2) for background.
-Our paper considers the following question: Given the standard experimental designs (e.g., product-level or customer-level randomization) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments?
+Our paper considers the following question: Given the standard experimental designs (e.g., listing-side randomization, LR, or customer-side randomization, CR) and estimators (i.e., difference-in-means) that are employed by platforms, what are the consequences of interference bias in pricing experiments?
 There are several key features of our approach.
@@
 Furthermore, our analysis focuses on differential changes in continuous variables such as prices, enabling us to leverage calculus to derive straightforward and intuitive expressions.
-(In practical settings, we will think of “small” changes around the control value.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
-We first introduce a succinct yet general model of platform experimentation in which a firm wants to understand the effect of a continuous variable on a metric of interest.
-The model is general in that can it encode different market arrangements and experimental designs.
+(In practical settings, this corresponds to “small” changes around the control price.) Finally, unlike many preceding studies on online platform experimentation, we focus on profits as an outcome metric, in addition to bookings or transaction volume.
+We first introduce in Section 3 a succinct yet general model of platform experimentation in which a firm wants to understand the impact of modifications to a continuous variable, on a metric of interest.
+The model is general, in the sense that it can represent various market arrangements and experimental designs.
 By leveraging differential calculus, we provide several structural results regarding the performance of canonical estimators and designs.
-Notably, we show that the bias corresponds to the cross-effects of the control and treatment metrics with respect to marginal changes on the other group’s variable.
+Notably, we show that the bias corresponds to the cross-effects of the control and treatment group metrics, with respect to marginal changes in the other group’s variable.
 These cross-effects are precisely an indicator of interference, providing an intuitive structural characterization of bias.
-Second, we specialize our results to a canonical revenue management pricing problem.
-By using our bias characterizations, we show that experiments can obtain the wrong sign when the metric of interest is profit (or revenue).
-In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive).
-(Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate obtains the correct sign.) We express the conditions under which an estimate with incorrect sign can be obtained with straightforward expressions that involve price elasticities.
-[^1] The expressions are reminiscent of elasticity conditions in Blake and Coey, 2014a .
-Our results have important implications for price optimization based on experiments: incorrectly following the experiment’s suggestion may lead one to move in the wrong direction and decrease profits relative to the status-quo!
-We take advantage of our structural characterization of bias to construct a novel debiasing method for the canonical A/B experiment.
-In particular, we show that if the experiment equally allocates units between treatment and control, then it is possible to fully debias the canonical estimator, by taking advantage of symmetry in the structure of bias.
+Second, in Sections 4 and 5, we specialize our results to a canonical revenue management pricing problem.
+By using our bias characterizations, we show that the bias can be severe enough that experimenters can obtain the wrong sign when the metric of interest is profit (or revenue).
+In particular, we show that when treatment involves raising prices (resp., lowering prices), the global treatment effect may be positive (resp., negative), but due to interference, the estimator can be negative (resp., positive) in expectation.
+Interestingly, when treatment involves raising prices (resp., lowering prices), and the global treatment effect is negative (resp., positive), the experimental estimate always obtains the correct sign.
+We express the conditions under which an estimate with the wrong sign can be obtained with straightforward expressions that involve price elasticities.
+[^2] The expressions are reminiscent of elasticity conditions in Blake and Coey, (2014), though their analysis differs from ours: they focus on estimating the relative bias in terms of supply and demand elasticities, and carry out a local change with respect to a multiplicative factor by which treatment demand changes relative to control.
+By contrast, our estimand is local with respect to the magnitude of the price change, and of course, a significant emphasis of our study is the possibility of mismatched signs between the true $GTE$ and the estimated $GTE$, which they do not study.
+In the appendix, we also show that these results generalize to a setting where a platform with heterogeneous goods for sale sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces.
+Our results have important implications for price optimization based on experiments: following the suggestion of a canonical experiment may lead one to move in the wrong direction and decrease profits relative to the status-quo!
+Taking advantage of our structural characterization of bias, in Section 6, we construct a novel debiasing method for canonical A/B experiments.
+In particular, we show that if the experimenter equally allocates units between treatment and control, it is possible to fully debias the canonical estimator by taking advantage of symmetry in the structure of bias.
 Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below).
-This is a stark contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require much more complicated experimental designs and/or estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
-To obtain additional insight, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022).
-We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs.
-For example, for a listing-side experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low.
-We also investigate numerically how bias depends on both market balance and the baseline control price.
-We conclude the paper with an alternative specialization of our revenue management setting to a model where the firm sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces.
-In this model we allow the possibility of a heterogeneous collection of goods for sale, with an unknown (joint) demand function.
-We show that all our previous results extend to this setting as well.
+This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
+To obtain additional insight, in Section 7, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022).
+We show how the biases of LR and CR experiments depend on model parameters, with a particular focus on regimes in which a “change of sign” may occur.
+For example, for an LR experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is low relative to supply.
+We also investigate numerically how bias depends on both market balance and the baseline control price, and show how canonical A/B methods for price experimentation may result in significant profit underperformance.
+In Section 8, we conclude the paper by calibrating our two-sided market model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S.
+cities.
+This exercise simplifies the Airbnb marketplace in many dimensions, but we believe it is useful for offering insights into the prevalence of estimators with the wrong sign in real-world markets.
+In our investigation, we find that such estimators are plausible in the Airbnb context.
+Notably, all markets in the study yielded estimators with the wrong sign when we simulated LR experiments, which would be a typical approach for testing pricing changes in practice.
+Although this analysis is based on a simplified market model, the results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners.
+We conclude in Section 9.
 2 Related Work
@@
 More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015).
-Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice.
-Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), and message passing (shirani2023causal).
-A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023).
-Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects).
-Roemheld and Rao, (2024) uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments.
-More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods.
+Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014, Fradkin, 2019, Liu et al., 2020), showing its importance in practice.
+Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
+A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023).
+Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation.
+More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects).
+Roemheld and Rao, (2024) uses an empirical study to suggest that false positives may be inflated due to interference in pricing experiments.
+More similar to our findings, Deng and Le, (2023) also notes the possibility of the sign of the treatment effect estimator being incorrect; in their setting this is due to an effect they call lead-day bias in markets with expiring goods.
 In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign.
-Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper – those where a continuous variable (such as price) is incrementally varied – as marginal policy changes, and develops a framework to study them.
-The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “naive” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments.
-In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price in our present paper.
-Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is more focused on design and estimation for general estimands in market equilibrium.
+Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper—those where a continuous variable (such as price) is incrementally varied—as marginal policy changes, and develops a framework to study them.
+The closest of these works to ours is Munro et al., (2021), which studies general estimands in an economic price equilibrium setting, and decomposes the treatment effect into direct and indirect effects; the former is obtained via “canonical” estimators, while the latter corresponds to our bias, and the authors show how it can be corrected using local price experiments.
+In that work, the direct and indirect effects are obtained as marginal changes with respect to the treatment assignment probability (inspired by work in Hu et al., (2022)), rather than the price as in our present paper.
+Further, our emphasis in applying our results is on structural insight regarding when changes of sign may be obtained for objectives such as profits or revenue, while their paper is focused on design and estimation for general estimands in market equilibrium.
 (An earlier version of Munro et al., (2021) considers treatment changes in a continuous variable, with an emphasis on relating the resulting differential treatment effect to direct and indirect effects.)
@@
 The firm is interested in optimizing the performance metric by adjusting the decision variable $x$.
-In particular, the firm is interested in the difference in the metric when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, relative to a status-quo “control” value, denoted by $x_{0}$.
+In particular, the firm is interested in the change in the target metric, when the decision variable takes on a new “treatment” value, denoted by $x_{1}$, moving away from the status-quo “control” value, denoted by $x_{0}$.
 This difference, $T(x_{1})-T(x_{0})$, corresponds to the global treatment effect ($GTE$).
-We define a local version of $GTE$ considering that the treatment value of the decision variable is “close” to the control value:
-$GTE(x_{0})$ ::=:= $T^{\prime}(x_{0})\ .$ (1)
+We define a local version of the $GTE$ as the derivative of the target metric, representing the case in which the treatment value of the decision variable is “close” to the control value:
+$GTE(x_{0})$ $:=T^{\prime}(x_{0})\ .$ (1)
 $GTE(x_{0})$ is our estimand.
@@
 Our work focuses on settings where the function $T$ may be unknown to the firm or platform.
-Therefore, an experiment is conducted to determine the direction in which the decision variable should be modified to optimize the performance metric.
+Therefore, an experiment is conducted to determine the direction in which the decision variable should be adjusted in order to improve the performance metric (a form of policy improvement).
 We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control.
-We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated in each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
-It is worth noting that in the settings we are considering, randomization can occur at either side of the market, i.e., customers or products (i.e., listings in the case of a platform) can be randomized.
-In a customer-side experiment, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition.
-On the other hand, in a product (listing)-side experiment, a customer is exposed to a mix of treatment and control products (or listings).
+We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated to each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
+It is worth noting that in the settings we are considering, randomization can occur on either side of the market, i.e., customers or products (or listings).
+In a customer-side randomization experiment (CR), a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition.
+On the other hand, in a listing-side randomization experiment (LR), a customer is exposed to a mix of treatment and control products (or listings).
 As before, we denote by $x_{0}$ the baseline (control) value of the decision variable and by $x_{1}$ the treatment value.
-We assume that the performance metric under the treatment and control conditions are given by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$, which we refer to as the control and treatment performance metrics, respectively.
-The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment value are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units is randomized into treatment.
+We denote the in-experiment performance metrics for control and treatment goods by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$ respectively.
+The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment values are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units are randomized into treatment.
 The nature of these functions will become more clear in the context of the applications described below.
+Notably, this notation abstracts away from the specific experimental design (whether CR or LR).
 We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
-The canonical experimental designs at each side of the market have an associated natural “naive” local estimator.
+The canonical experimental designs at each side of the market have an associated natural “canonical” local estimator.
 To define this estimator, consider a sequence of experiments for which the treatment value of the decision variable approaches the control value.
 Then, we compute the relative difference between the performance metrics appropriately normalized by the treatment and control fractions, respectively, which corresponds to the Horvitz-Thompson estimator, or equivalently in our formal mean-field setting, the difference-in-means estimator.
-More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, for which the treatment approaches control: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$.
-We define a local version of the naive estimator as:
-$\widehat{GTE}(x_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (2)
+More formally, we consider a sequence of experiments indexed by $n\in\mathbb{Z}_{+}$, in which the treatment value approaches the control value: $\lim_{n\to\infty}x_{1}^{n}=x_{0}$.
+We define a local version of the canonical estimator as:
+$\widehat{GTE}(x_{0},q)$ $:=\lim_{n\to\infty}\frac{\frac{T_{1}(x_{0},x_{1}^{n},q)}{q}-\frac{T_{0}(x_{0},x_{1}^{n},q)}{(1-q)}}{x_{1}^{n}-x_{0}}\ .$ (2)
 We make the following assumption that is kept throughout the paper:
@@
 Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized.
-Note that experimental estimates in practice typically implicitly rely on the assumption of constant returns to extrapolate estimates for control and treatment groups to the entire market.
+Note that in practice, experimental estimates typically implicitly rely on the assumption of constant returns in order to extrapolate estimates for control and treatment groups to the entire market.
 For the marketplace models and metrics we study later, this assumption will be satisfied.
 Now, we obtain:
-$\widehat{GTE}(x_{0},q)$ $=$ $\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$ (3)
+$\widehat{GTE}(x_{0},q)$ $=\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{1}(x_{0},x_{0},q)/q-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
 $=\lim_{n\to\infty}\frac{T_{1}(x_{0},x_{1}^{n},q)/q-T_{1}(x_{0},x_{0},q)/q+T_{0}(x_{0},x_{0},q)/(1-q)-T_{0}(x_{0},x_{1}^{n},q)/(1-q)}{x_{1}^{n}-x_{0}}$
-$=$ $T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q),$
+$=T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\ .$ (3)
 The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.2.
 The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
-Finally, we define the estimation bias of this local estimator as:
-$Bias(x_{0},q)$ ::=:= $GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
-$=$ $T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)$ (5)
+Finally, we define the estimation bias of this local estimator as: [^3] Note that in statistics, bias is often defined as the estimator minus the estimand.
+However, we make the preceding choice because (as we will see) the resulting sign of $Bias$ is typically nonnegative, making it easier to obtain practical interpretations of our results.
+$Bias(x_{0},q)$ $:=GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
+$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .$ (5)
 Studying bias in this local regime proves fruitful, as it leads to the following simple characterization.
@@
 The bias is given by the following expression:
-$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q$
+$Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q\ .$
 Proof 3.2
@@
 $Bias(x_{0},q)=T_{0y}(x_{0},x_{0},q)/(1-q)+T_{1x}(x_{0},x_{0},q)/q\ .$
+$\Box$
 The result provides an intuitive characterization of bias.
-Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable on the other group changes marginally.
+Specifically, $T_{0y}(x_{0},x_{0},q)$ and $T_{1x}(x_{0},x_{0},q)$ correspond to the relative changes in performance for the control and treatment groups, respectively, when the decision variable of the other group changes marginally.
 The cross-partial derivatives represent the effect on the performance metric across experimental groups, making them an intuitive indicator of interference.
-As interference increases, bias correspondingly intensifies.
+As expected, as this interference effect increases, bias correspondingly increases.
 4 Revenue Management Setting
-We examine a traditional revenue management scenario involving monopoly pricing.
+In this section we examine a traditional revenue management setting with monopoly pricing.
 A firm or platform offers products that, at least based on observable characteristics, are deemed homogeneous and are thus priced uniformly.
-The firm faces an unknown to them demand function given by $D(p)$ as a function of the price $p$.
+The firm faces an (unknown to them) demand function given by $D(p)$ as a function of the price $p$.
 The firm has a constant marginal production cost, $c\geq 0$.
@@
 At this level of abstraction, the demand function can describe a scenario with either an infinite supply or a limited supply, where a rationing mechanism determines the consumption.
-In a large limit model, even with supply constraints we expect the demand function to be differentiable.
-To provide a concrete example with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price.
-That is, $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$ almost surely, where $\{x_{i}(p):i=1,...,n\}$ are i.i.d.
-random variables equal to one if customer $i$ buys the product at a given price $p$, and zero otherwise.
+In the limit of a large market, even with supply constraints, we expect the demand function to be differentiable.
+To provide a concrete example of a limiting regime with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price.
+That is, suppose $x_{i}(p)\in\{0,1\},i=1,2,\ldots$ are i.i.d.
+Bernoulli random variables with mean $0\leq D(p)\leq 1$; we interpret $x_{i}(p)$ as the purchasing decision of customer $i$ at price $p$.
+Then trivially by the law of large numbers $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$.
 4.1 Estimands
-We will consider two performance metrics of typical interest: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero).
-The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand) and subsequently affect profit.
+We will consider two performance metrics: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero).
+The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand), and subsequently affect profit.
 To achieve this goal, the firm is interested in estimating the difference in demand or profits generated when products are priced at a new “treatment” price relative to the demand or profits generated when products are priced at the status-quo “control” price.
@@
 Following equation (3.1), we present our estimands. These are the local versions of the global treatment effects associated to demand and profits:
-$GTE_{D}(p)$ ::=:= $D^{\prime}(p)$ (6)
-$GTE_{\pi}(p)$ ::=:= $\pi^{\prime}(p)$ (7)
-$=$ $D(p)+(p-c)D^{\prime}(p)$
+$GTE_{D}(p)$ $:=D^{\prime}(p)\ .$ (6)
+$GTE_{\pi}(p)$ $:=\pi^{\prime}(p)$
+$=D(p)+(p-c)D^{\prime}(p)\ .$ (7)
 Note that $GTE_{\pi}(p)$ is the marginal profit of the firm at price $p$.
@@
 Given that the demand function is unknown, we assume the firm uses price experimentation to estimate the $GTE$s of interest.
-As mentioned above, canonical A/B-tests can randomize at either side of the market, and our model does not preclude either possibility.
+As mentioned above, canonical A/B-tests can randomize on either side of the market, and our model does not preclude either possibility.
 That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
@@
 $D_{0}(p_{0},p_{1},q)=\frac{Q_{0}(p_{0},p_{1},q)}{1-q};\quad D_{1}(p_{0},p_{1},q)=\frac{Q_{1}(p_{0},p_{1},q)}{q},$
-which correspond to the demand for control and treated subjects scaled by the control and treatment fractions, respectively.
-Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions.
+which represent the demand for control and treated units scaled by the control and treatment fractions, respectively.
+Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions, respectively.
 Again, at this level of abstraction these demand functions could, for example, incorporate supply constraints and rationing mechanisms.
-Recall that in a customer-side experiment treatment and control customers will observe all products priced at the treatment and control prices, respectively, with the proportion of customers assigned to the treatment group being $q$.
-In contrast, in a product-side experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
+Recall that in a CR experiment treatment and control customers will observe all products priced at the treatment and control prices respectively, with the proportion of customers assigned to the treatment group being $q$.
+In contrast, in a LR experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
 We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper.
@@
 Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous.
-Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
-We introduce the following additional assumption on experimental demand functions that implies Assumption 3.2. {assumption} The following holds:
+Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways, commonly assumed in pricing settings: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
+We introduce the following additional assumption on the experimental demand functions which implies Assumption 3.2. {assumption} The following statements all hold:
   1. 1.
@@
     $D_{0}(p,p,q)=D_{1}(p,p,q)=D(p)$, for all $p\in\mathbb{R}_{+}$ and $10<q<$ 1.
-The first two assumptions are natural as they imply that control and treatment products are (imperfect) substitutes.
+The first two statements in Assumption 4.2 are natural as they imply that control and treatment products are (imperfect) substitutes.
 The third assumption states that if the control and treatment prices are the same, then both the scaled control and treatment demand functions are equivalent to the market demand function; this assumption is implied by constant returns to scale.
@@
 We finish this subsection by defining the experimental estimators. To do so, we define the experimental control and treatment profit functions, respectively:
-$\pi_{0}(p_{0},p_{1},q)$ ::=:= $(p_{0}-c)D_{0}(p_{0},p_{1},q)$
-$\pi_{1}(p_{0},p_{1},q)$ ::=:= $(p_{1}-c)D_{1}(p_{0},p_{1},q).$
-Following definition (2) we define naive estimators for demand and profits associated to the canonical designs:
-$\widehat{GTE}_{D}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
-$\widehat{GTE}_{\pi}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
-Note that analogs of these naive estimators are frequently used in practice in estimating global treatment effects from experimental data.
+$\pi_{0}(p_{0},p_{1},q)$ $:=(p_{0}-c)D_{0}(p_{0},p_{1},q)\ ,$
+$\pi_{1}(p_{0},p_{1},q)$ $:=(p_{1}-c)D_{1}(p_{0},p_{1},q)\ .$
+Following Definition (2) we define canonical estimators for demand and profit associated with the canonical experimental designs:
+$\widehat{GTE}_{D}(p_{0},q)$ $:=\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}\ ,$
+$\widehat{GTE}_{\pi}(p_{0},q)$ $:=\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}\ .$
+We note that analogs of these canonical estimators are frequently used in practice in estimating global treatment effects from experimental data; in particular, in our continuum model, these estimators naturally correspond to difference-in-means or Horvitz-Thompson type estimators, scaled by the price difference to estimate the target estimand (the derivative of demand or profit at the control price).
 Now, by equation (3), we obtain:
 $\widehat{GTE}_{D}(p_{0},q)$ $=D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\ ,$ (8)
-$\widehat{GTE}_{\pi}(p_{0},q)$ $=$ $\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$ (9)
-$=$ $D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$
+$\widehat{GTE}_{\pi}(p_{0},q)$ $=\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$
+$=D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$ (9)
 where we used Assumption 4.2 and the last equation follows by Assumption 4.2.3.
-5 Biases: Analysis and Results
-In this section we analyze biases associated to the naive estimators. Following definition (4) we define the estimation bias for demand and profits associated to the canonical designs:
-$Bias_{D}(p_{0},q)$ ::=:= $GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)$
-$Bias_{\pi}(p_{0},q)$ ::=:= $GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)$
+5 Biases: Analysis and Change of Sign
+In this section we analyze the biases associated with the canonical estimators introduced above. Following Definition (4), we define the estimation biases for demand and profit as:
+$Bias_{D}(p_{0},q)$ $:=GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)\ ,$
+$Bias_{\pi}(p_{0},q)$ $:=GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)\ .$
 First, note that by Proposition 3.1 we obtain the following bias characterizations:
-$Bias_{D}(p_{0},q)$ $=$ $D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)$ (10)
-$Bias_{\pi}(p_{0},q)$ $=$ $\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$ (11)
-$=$ $(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$
-$=$ $(p_{0}-c)Bias_{D}(p_{0},q)$ (12)
-As discussed in Proposition 3.1, bias increases with increasing cross-effects between treatment and control groups.
-More specifically, the bias is influenced by the degree to which the control (or treatment) group detracts demand from its counterpart in response to a marginal price change in the latter group.
-Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, so we obtain that the naive estimators always (weakly) underestimate $GTE$:
+$Bias_{D}(p_{0},q)$ $=D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\ ,$ (10)
+$Bias_{\pi}(p_{0},q)$ $=\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$
+$=(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$ (11)
+$=(p_{0}-c)Bias_{D}(p_{0},q)\ .$ (12)
+As discussed in Proposition 3.1, bias increases as interference effects between treatment and control groups increase.
+More specifically, the bias is influenced by the degree to which the control (or treatment) group cannibalizes demand from its counterpart in response to a marginal price change in the latter group.
+Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, meaning the canonical estimators always (weakly) underestimate the $GTE$:
 Proposition 5.1
 We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
-The result follows similar cannibalization arguments developed in previous work (Johari et al., 2022).
-To provide intuition consider a listing-side experiment in which treatment price is larger than control price.
-In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from the treatment to the control group.
+This result follows similar cannibalization arguments developed in previous work (Johari et al., 2022).
+To provide intuition, consider an LR experiment in which the treatment price is higher than control the price.
+In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from treatment units to control units.
 Similarly, experimental control demand overestimates global control demand.
-Now, let’s consider a customer-side experiment where there are supply constraints, so customers compete for limited products.
-In such cases, the control customers may “cannibalize” supply from the treatment group, given the lower control price, leading again to an underestimation of the global treatment effect.
-Note that in a customer-side experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
+Instead, let’s consider a CR experiment where there are supply constraints, so customers compete for limited products.
+In such cases, the control customers may “cannibalize” supply from the treatment group, because of their lower control price, causing an underestimation of the global treatment effect again.
+Note that in a CR experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
 It is important to observe that because $D^{\prime}(p)<0$ by Assumption 4, $GTE_{D}(p_{0})<0$, for all $p$.
-Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the naive estimator will have the same (nonpositive) sign.
+Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the canonical estimator will have the same (nonpositive) sign.
 However, there is a more intriguing result for profits.
-Suppose that the price $p_{1}$ is greater than $p_{0}$.
+Suppose that the treatment price $p_{1}$ is greater than the control price $p_{0}$.
 In this scenario, even if the demand $D(p_{1})$ at $p_{1}$ is lower than the demand $D(p_{0})$ at $p_{0}$, the increase in price may offset the decrease in demand, leading to a positive true global treatment effect in profits.
-However, because the naive estimator underestimates the treatment effect, it could be negative.
-Next, we indeed provide conditions under which the naive estimator not only underestimates the global treatment effect, but also obtains the wrong sign.
+But, because the canonical estimator underestimates the treatment effect, the profit GTE estimator could be negative.
+Next, we will provide conditions under which the canonical estimator not only underestimates the global treatment effect, but actually obtains the wrong sign.
 By definition, $GTE_{\pi}(p_{0})\geq 0$ if and only if $\pi^{\prime}(p_{0})\geq 0$, which after some simple manipulation yields the following standard “markup” inequality:
@@
 For the case $c=0$, condition (a) naturally states that marginal revenue is increasing at $p_{0}$, or equivalently that the demand is relatively inelastic with respect to price.
-Now, consider the naive estimator.
+Now, consider the canonical estimator.
 Note that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ is equivalent to $Bias_{\pi}(p_{0},q)\geq{GTE}_{\pi}(p_{0})$.
-That is, provided that condition $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect.
+That is, provided that $GTE_{\pi}(p_{0})\geq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect.
 Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
 […]
-Note that condition (b) is a “modified markup” inequality. We have the following result:
+Note that condition (b) is a “modified markup” inequality. We have the following result.
 Proposition 5.2
@@
 $\frac{p_{0}-c}{p_{0}}\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\leq\frac{p_{0}-c}{p_{0}}D^{\prime}(p_{0})\ ,$
-and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4.
-The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold, that is there are appropriate lower and upper bounds on the inverse of the price elasticity.
+and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4. $\Box$
+The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold.
+This implies lower and upper bounds on the inverse of the price elasticity of demand.
 Note that we use weak inequalities in conditions (a) and (b) for a change of sign; strictly speaking of course, a change of sign can only occur if at least one of the two inequalities is strict.
@@
 [^5] In our two-sided market model, introduced in section 7, the scenario $\widehat{GTE}_{\pi}(p_{0},q)=GTE_{\pi}(p_{0})=0$ is impossible, meaning this edge case does not exist.
-To provide intuition, note that by using equation (8) we can define the experimental price elasticity of demand as:
+To provide intuition, note that by using equation (8) we can define the “experimental” price elasticity of demand as:
 $\widehat{e_{p}}(p_{0},q)=\frac{p_{0}(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q))}{D(p_{0})}.$ (15)
@@
 which can be understood as the experimental analogue of a markup inequality.
-It states that at the experimental elasticity, the current price is above the optimal monopoly price (under an appropriate concavity assumption).
-Hence, as expected, a change of sign occurs when the naive estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices).
-The result has important consequences for experiment-driven price optimization: the prescription implied by the experiment (reduce price) is the opposite to the one implied by the $GTE$ (increase price).
-Hence, if one moves in the direction suggested by the experiment, profits decrease relative to the status-quo.
+It states that the experimental elasticity suggests the current price is above the optimal monopoly price (under an appropriate concavity assumption).
+Hence, as expected, a change of sign occurs when the canonical estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices).
+This result has important consequences for experiment-driven price optimization: the prescription implied an the experiment (reduce price) may be the opposite of the one implied by the $GTE$ (increase price).
+Therefore, if a firm moves in the direction suggested by the experiment, profits have the potential to decrease relative to the status-quo.
 To provide further intuition, let us assume that condition (a) holds and that profits are quasiconcave, so that the firm is underpricing relative to the optimal monopoly price.
@@
 A change of sign happens as the two cannibalization effects just mentioned become larger.
-We conclude with a remark regarding our estimators, our findings on changes of sign, and guidance for practitioners.
-Note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit function, respectively, as is natural given our estimands.
-In practice, a platform will typically directly use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices.
+We conclude with two remarks regarding our estimators, our findings on changes of sign, and guidance for practitioners.
+First, note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit functions, respectively, as is natural given our estimands.
+In practice, a platform may use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices.
 This can have some bearing in interpreting our results, depending on the sign of the difference between the treatment and control price.
 Concretely, suppose that treatment involves raising the price relative to control.
-Then our results show that a change of sign can only occur if the true global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment.
-In other words, in this case, the platform risks failing to move prices in a profitable direction.
+Then our results show that a change of sign can only occur if the unnormalized global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment.
 On the other hand, suppose the treatment involves lowering the price relative to control.
-Then our results show that a change of sign can only occur if the true global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment.
-In other words, in this case, the platform risks moving prices in a potentially unprofitable direction.
+Then our results show that a change of sign can only occur if the unnormalized global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment.
+In either case, the platform risks either failing to move prices in a profitable direction, or worse, moving them in an unprofitable direction.
+Second, we note that in our case, we have studied a model where the platform changes the price directly.
+In practice, another relevant platform parameter that affects profits can be a platform-wide fee.
+Specifically, in Appendix 12 , we present a model where a platform offers a multitude of heterogeneous goods (potentially at different prices), and earns profit by charging a percentage fee on each transaction.
+This is a common monetization model for online marketplaces.
+Viewing the fee as the variable of interest for the platform’s profit optimization, we obtain analogous results to those in this section, including the bias characterization, the possibility of a change of sign, and interpretation of the conditions for a change of sign in terms of appropriate derivatives .
+This extension suggests our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.
 6 Debiasing Canonical Estimators
-In this section, we take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that applies as long as the platform has equally allocated units between treatment and control, i.e., $q=1/2$.
-Recall (10) (11) provide the expressions for bias in terms of derivatives of demand.
-We now show how we can use observables to estimate this bias.
+We take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$.
+The key to this approach, and a necessary step for it to be applied, is for the experimenter to equally allocate units between treatment and control, i.e., setting $q=1/2$.
+Recall that equations (10) and (11) quantify bias in terms of the derivatives of demand.
+We will now show how we can use observables to estimate this bias.
 First, we show that $D_{0y}(p_{0},p_{0},q)$ can be computed (approximately) using observed data.
@@
 By Assumption 6 we have that $D_{0}(p_{0},p_{0},q)=D(p_{0})$.
-Since we assume that the firm was pricing at $p_{0}$ before the experiment began, they have observed $D(p_{0})$ prior to the experiment.
+To make progress, we make the assumption that since the firm was pricing at $p_{0}$ before the experiment began, the pre-experiment observation of demand is a reasonable proxy for the control demand $D(p_{0})$.
 The quantity $D_{0}(p_{0},p_{1},q)$ is observed as well: it represents the control demand during the experiment.
@@
 Taking the derivative of both sides of this equality with respect to $p_{1}$, we obtain $D_{0y}(p_{0},p_{1},1/2)=D_{1x}(p_{1},p_{0},1/2)$, so that when $p_{1}=p_{0}$, there holds $D_{0y}(p_{0},p_{0},1/2)=D_{1x}(p_{0},p_{0},1/2)$.
-We have shown the following result.
+We have thus shown the following result.
 Proposition 6.1
@@
 $GTE_{\pi}$ $=\widehat{GTE}_{\pi}(p_{0},1/2)+2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (19)
-Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables, then the right hand sides of (16) and (17) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
-We note several limitations of this method.
+Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables (as described above), then the right hand sides of (18) and (19) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
+We note three limitations of this method.
 First, of course, the symmetry property being exploited only applies when the market is equally allocated to treatment and control.
-Second, the assumption that $D(p_{0})$ is observed relies on market stationarity; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment.
-Finally, estimation of $D_{0y}$ uses only half of the experimental data, namely, only the observed control demand, which may be a source of sample inefficiency (i.e., increased variance).
+Second, the method is not sample efficient: only half the experimental samples, those from the control group, are used.
+Finally, our view that the platform can accurately estimate $D(p_{0})$ depends on the market remaining reasonably stationary during the experiment; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment.
 7 A Two-Sided Market Model
-In this section, we illustrate the application of our results of the general revenue management setting introduced in Section 4 to a two-sided market model developed in Johari et al., (2022).
+In this section, we illustrate the application of our results from the general revenue management setting (introduced in Section 4) to the two-sided market model developed in Johari et al., (2022).
 That paper develops a stochastic dynamic market model with supply constraints in which customers book listings for a finite period of time.
@@
 We begin by recapping the mean-field model formulation in Johari et al., (2022), adapted to our setting with price experiments.
-There is a two-sided platform evolving in continuous time with listings on the supply-side and customers on the demand side.
+There is a two-sided platform evolving in continuous time with listings on the supply side and customers on the demand side.
 There is a continuum of listings of mass $\rho$ that are homogenous with respect to their observable characteristics.
@@
 […]
-$q(s)=\frac{sv(p)}{\varepsilon+sv(p)}.$
+$\frac{sv(p)}{\varepsilon+sv(p)}.$
 Finally, a listing remains occupied for an expected time $\frac{1}{\tau}$ when it gets booked; in our analysis we normalize $\tau$ to be 1 without loss of generality.
 Note that in our model, the parameter $\lambda$ mediates market balance: when $\lambda\rightarrow 0$ the market becomes demand constrained, when $\lambda\rightarrow\infty$ the market becomes supply constrained.
-Note that the general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers. In the model we consider in this section, we make both consumers and listings homogeneous.
+The general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers.
+To simplify the analysis and exposition, in the model we consider in this section, we make both consumers and listings homogeneous.
 As in Johari et al., (2022) we focus on the steady-state of the system.
@@
 7.3 Experimental Designs
-In this section we analyze two canonical experimental designs in platforms: randomizing listings or randomizing customers.
-7.3.1 Listing-Side Randomization
-We assume that randomization occurs on the listing side, where a fraction $q$ of the listings is assigned to treatment.
+In this section we analyze two canonical experimental designs in platforms: randomization across listings (LR experimentation) and randomization across customers (CR experimentation).
+7.3.1 LR Experimentation
+We assume that randomization occurs on the listing side, where a fraction $q$ of the listings are assigned to treatment.
 Hence, a customer is exposed to a mix of treatment and control listings.
@@
 Note that the bias is independent of the treatment fraction $q$.
-7.3.2 Customer-Side Randomization
-In this case we randomize on the customer side, with a fraction $q$ of customers randomized into treatment.
+7.3.2 CR Experimentation
+In this setting, we randomize over customers, with a fraction $q$ of customers being placed into treatment, and the remainder in control.
 In this case, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition.
@@
 $\rho-s^{*}(\lambda,p_{0},p_{1},q)=\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (25)
-In this setting we have the following demand functions:
+Observe that the two terms on the right hand side above are rescaled versions of the control and treatment demand functions:
 $D^{CR}_{0}(\lambda,p_{0},p_{1},q)$ $=\lambda\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})},$ (26)
@@
 7.4 Properties of Biases with Market Balance
-In this section we study structural properties of biases with a particular focus on extreme regimes of market balance. The analysis in this subsection closely follows Johari et al., (2022).
-[…]
-As the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines resulting in less interference in a CR experiment.
-In the limit, there is no competition between customers and we expect the bias in the CR experiment to be zero.
-On the other hand, competition between listings increases resulting in more interference in a LR experiment.
-In the limit, we expect the bias in the LR experiment to be non-zero.
-Conversely as the market becomes supply constrained ($\lambda\rightarrow\infty$), competition between listings declines resulting in less interference in a LR experiment.
-In the limit, there is no competition between listings and we expect the bias in the LR experiment to be zero.
-Competition between customers increases resulting in more interference in a CR experiment.
-In the limit, we expect the bias in the CR experiment to be non-zero.
-The following propositions formalize the intuition about the behavior of the bias in both extremes of market balance.
-Proposition 7.1
-For all $p\in[c,+\infty)$, we have that:
-$\lim_{\lambda\to 0}GTE_{D}(\lambda,p)/\lambda=\frac{\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}};\ \ \lim_{\lambda\to 0}GTE_{\pi}(\lambda,p)/\lambda=\frac{\rho v(p)}{\varepsilon+\rho v(p)}+\frac{(p-c)\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}}.$
-We have that $\lim_{\lambda\to 0}Bias_{\pi}^{CR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{CR}(\lambda,p)/\lambda=0$, while
-$\lim_{\lambda\to 0}Bias_{\pi}^{LR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{LR}(\lambda,p)/\lambda=-\frac{(p-c)\rho^{2}v(p)v^{\prime}(p)}{(\varepsilon+\rho v(p))^{2}}.$
-Proposition 7.2
-For all $p\in[c,+\infty)$, we have that: $\lim_{\lambda\to\infty}GTE_{D}(\lambda,p)=0$ and that:
-$\lim_{\lambda\to\infty}GTE_{\pi}(\lambda,p)=\rho$. We have $\lim_{\lambda\to\infty}Bias_{\pi}^{LR}(\lambda,q)=(p-c)\lim_{\lambda\to\infty}Bias_{D}^{LR}(\lambda,q)=0$, while
-$\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=(p-c)\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=-\frac{\rho(p-c)v^{\prime}(p)}{v(p)}.$
-In the limit where $\lambda\rightarrow 0$, Proposition 7.1 suggests that the CR experiment is unbiased while the LR experiment is biased.
-Similarly, in the limit where $\lambda\rightarrow\infty$, Proposition 7.2 suggests that the CR experiment is biased while the LR experiment is unbiased.
-Proofs of both propositions are presented in Appendix 11.
-The proofs are similar to the analysis in Johari et al., (2022) but we also leverage our characterization of biases based on cross-partial derivatives.
-In the next section we will further study the extent of these biases to characterize change of sign regimes.
-We also provide some figures to complement our limit results.
-In Figure 1.
-As $\lambda$ increases, the market becomes more supply-constrained and bias decreases as competition between listings dissipates.
-Figure 2 shows the normalized bias of profits for CR, and this dependence on $\lambda$ is naturally reversed.
-Both figures also show interesting behavior of biases with respect to the baseline price $p$.
-As the price $p$ changes, the sensitivity of demand to price changes as well; in particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit.
+Johari et al., (2022) studies the bias of CR and LR experiments as the market becomes heavily demand constrained or supply constrained, respectively.
+Following the analysis of that paper, we obtain analogous insights, with proofs similar to Johari et al.
+(2022) but also leveraging our characterization of biases through cross-partial derivatives.
+In particular, as the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines, resulting in less interference in a CR experiment; in the limit, the bias in a CR experiment approaches zero.
+On the other hand, in the same regime the competition between listings is high, resulting in more interference for an LR experiment, so in the limit the bias of an LR experiment remains positive.
+By contrast, as the market becomes more supply constrained ($\lambda\to\infty$), the opposite results hold: in this regime, competition between customers is stronger and competition between listings is weaker, so in the limit the bias of a CR experiment remains positive while the bias of a LR experiment approaches zero.
+We derive these results in Propositions 13.1 and 13.2 in Appendix 13.
+Below we provide two figures to complement our limit results.
+In Figures 1 and 2 we present bias normalized by the arrival rate $\lambda$.
+We normalize by $\lambda$ instead of the GTE because the GTE for profits converges to 0 around the optimal price.
+As $\lambda$ increases, the market becomes more supply-constrained and bias for a LR experiment decreases as competition between listings dissipates.
+Figure 2 shows the normalized bias of demand and profits in a CR experiment and this dependence on $\lambda$ is naturally reversed.
+Both figures also show interesting behavior of biases with respect to the baseline price $p$, a novel aspect of our setting relative to Johari et al., (2022).
+As the price $p$ changes, the sensitivity of demand to price changes as well.
+In particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit.
+Note in particular that LR experiments exhibit interesting behavior when prices are large: despite the fact that we nominally expect bias to increase when the market is demand constrained, and higher prices reduce demand in the market, nevertheless the bias drops when prices are high.
+This is because of the fact that elasticities change with price, and illustrates the nuanced nature of interference in price experiments.
 Figure 1: Illustration of $Bias^{LR}_{D}/\lambda$ and $Bias^{LR}_{\pi}/\lambda$ as a function of price ($p$) and market balance ($\lambda$).
@@
 […]
-An interesting immediate observation is that the regimes in which a change in sign occurs in an LR or CR experiment using the corresponding naive estimators are independent of the treatment fraction $q$, because the corresponding biases are independent of $q$.
-This has important implications for practice, as the treatment fraction is not a lever that can prevent the possibility that the sign of the treatment effect estimate is incorrect.
-Using the structure of the model we can provide the following insightful characterizations for the change of sign regions, as follows.
+Interestingly, observe that the bias terms derived in Section 7.4 do not depend on the treatment fraction $q$ for either LR or CR experiments.
+In other words, the treatment fraction is not a lever that can be used to prevent—or even reduce—the possibility of a change of sign in the treatment effect estimator.
+Using the structure of the model we can provide the following insightful characterizations for the change of sign regions.
 Theorem 7.1
-For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the listing-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{LR}_{\lambda}$.
+For all $\lambda\in(0,\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of an LR experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p\in[c,\infty)$ will be wrong if and only if $p\in P^{LR}_{\lambda}$.
 Theorem 7.2
-For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the customer-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{CR}_{\lambda}$.
-Theorems 7.3 and 7.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run.
+For all $\lambda\in(0,\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of a CR experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p\in[c,\infty)$ will be wrong if and only if $p\in P^{CR}_{\lambda}$.
+Theorems 7.1 and 7.2 state that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$, such that the sign of the canonical estimator will be wrong, regardless of whether an LR or CR experiment is run.
 The proofs are provided in Appendix 15.7 and repeatedly exploit the market balance equations for CR and LR respectively.
-The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$.
+The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}(\lambda)$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$, for each $\lambda$.
 In particular, this reveals that the sets $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not comprised of unrealistic prices: if a firm is currently pricing optimally or very slightly below optimally, the sign of their canonical estimator will be wrong, regardless of their experimental setup.
 It is not true, however, that the two experimental setups are equally fallible in all settings.
-Note that $\widehat{e_{p}}^{LR}$ is not equal to $\widehat{e_{p}}^{CR}$, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same, and behave quite differently based on market balance, as we illustrate through the following numerical study.
+Note that Condition (b) in 14 is different for LR and CR because the elasticities and demand sensitivities differ, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same.
+As expected based on the results of the previous section, they behave quite differently as a function of market balance, as we illustrate through the following numerical study.
 7.5.1 Numerical Results
 Figure (3) shows numerical results on the set of prices where a change of sign occurs.
-Figure 3: Illustration of the change of sign region (white) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$.
-The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in black.
-The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in grey.
-Consistent with Propositions 7.1 and 7.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment.
-At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest.
+Figure 3: Illustration of the change of sign region (purple) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$.
+The vertical axis is the (log of) the arrival rate, or market balance, paramter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted.
+The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red.
+The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
+Consistent with Propositions 13.1 and 13.2, in Figure 3 we observe that the change of sign region becomes very small in the demand-constrained regime for a CR experiment and in the supply constrained regime for an LR experiment.
+At intermediate market balance regimes, we can more immediately observe that the change of sign region corresponds to a single continuous interval on prices, as Theorems 7.1 and 7.2 suggest.
 We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected.
 We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
-8 Fee Model
-In this section we consider an environment in which a firm charges a fixed sales fee, as opposed to setting a price and incurring a marginal cost for each purchase.
-This represents the common situation–particularly in two-sided online marketplaces–where firms do not directly control the prices of the goods on their platforms, but instead collect a fixed fee for the sales of heterogeneous goods.
-8.1 Specifications
-Consider $N$ products, with a fixed vector of prices $\vec{p}\in\mathbb{R}^{N}$.
-The demand function $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$, takes in a vector of prices and outputs a vector in $\mathbb{R}^{N}$ that represents the quantity demanded for each good.
-We assume that the platform selects a single fee for all products $f\in\mathbb{R}^{+}$; this fee represents a percentage of all money transacted that the platform will collect.
-We assume that this fee is passed on entirely to the customer and that the platform’s marginal cost is zero.
-Therefore, we can write the platform’s profit $\Pi(f,\vec{p})$ as follows: [^10] Prices may be a function of the fee, since firms may pass-through fee changes.
-However, since experiments tend to be short-lived we ignore these longer term calibration effects.
-$\Pi(f,\vec{p})=f\vec{p}\cdot D((1+f)\vec{p}).$ (26)
-Let $D^{\prime}(\cdot)$ represent the matrix of demand derivatives in price. We can then write the platform’s profit derivative in the fee as:
-$\Pi_{f}(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p})+f\vec{p}\cdot\vec{p}\cdot D^{\prime}((1+f)\vec{p}).$ (27)
-Define $R(f,\vec{p})$ as the total money transacted in the platform when the fee is $f$ and the prices are $\vec{p}$. We can then write $R(f,\vec{p})$ as
-$R(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p}).$
-Let $R_{f}(f,\vec{p})$ represent the derivative of this quantity with respect to the fee. We can then rewrite (27) as:
-$\Pi_{f}(f,\vec{p})=R(f,\vec{p})+fR_{f}(f,\vec{p}).$ (28)
-And rewrite (26) as:
-$\Pi(f,\vec{p})=fR(f,\vec{p}).$
-{assumption}
-We assume that $\Pi_{f}$ exists and is well defined for all $f\in\mathbb{R}^{+}$ and all $\vec{p}\in\mathbb{R}^{N}$, or in other words, that $\Pi$ is differentiable over its domain.
-When running an A/B experiment, the platform will compare a control fee $f_{0}$ to a treatment fee $f_{1}$.
-To keep the presentation succinct, in this section we focus only on estimation of profits.
-This would cause a true global treatment effect (GTE) of $\Pi(f_{1},\vec{p})-\Pi(f_{0},\vec{p})$.
-Thus we define the (infinitesimal) $GTE$ at the control fee $f_{0}$ as:
-$GTE_{\Pi}(f_{0},\vec{p})=\Pi_{f}(f_{0},\vec{p}).$ (29)
-We assume that a fraction $q$ of all subjects are put in treatment and $1-q$ into control.
-Note that as in the monopoly pricing example, this randomization can occur on either the customer or product.
-For simplicity, in product side randomization, we assume that for each of the $n$ types of products, a fraction $q$ of each is in treatment treatment.
-Let $\Pi_{0}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from control subjects during the experiment, and $\Pi_{1}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from treatment subjects.
-Let $R_{0}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for control goods, and $R_{1}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for treatment goods.
-We then write $\Pi_{0}(f_{0},f_{1},\vec{p},q)=f_{0}R_{0}(f_{0},f_{1},\vec{p},q)$ and $\Pi_{1}(f_{0},f_{1},\vec{p},q)=f_{1}R_{1}(f_{0},f_{1},\vec{p},q)$.
-We make the following assumptions for these functions, analogous to Assumption 4.2.
-{assumption}
-The following holds:
+7.5.2 Gap from Optimality
+One important implication of Figure 3 is the impact that the change of sign region has on the net profits of firms who use canonical price experimentation methods.
+Consider the practical importance of the change of sign region’s borders.
+The right border of the change of sign region (where purple meets red), represents the optimal price for that particular market balance parameter $\lambda$.
+This is the point where $GTE_{\pi}=0$.
+The left border of the change of sign region (where blue meets purple) represents the final price a firm would reach, were it to repeatedly update it’s price in response to canonical A/B experimentation.
+This is the point where $\widehat{GTE}_{\pi}=0$, or equivalently, where $Bias_{\pi}=GTE_{\pi}$.
+Note that this price corresponds to the optimum suggested by repeated canonical A/B experiments.
+This is because, to the right of that price point, the expriment suggests decreasing the price, and to the left, to increase it.
+The difference between these two prices (the true optimal price and the experimentally optimized price) can correspond to significant losses in overall profit.
+In Figure 4, we see the percent of true optimal profit that a firm will achieve if it uses canonical LR (left) or CR (right) experimentation, for a range of market balance scenarios.
+What we find is that the profit achieved from optimization via experimentation can be far below the optimal profit, with the impact worsening in highly demand-constrained markets (for LR experiments) or demand-constrained markets (for CR experiments).
+Even when markets are reasonably balanced (i.e., $\lambda\approx 1$), both LR and CR experimentation will lead to a profit that is more than 4% lower than optimal.
+Overall, these findings suggest that firms who use canonical A/B methods for price experimentation might be operating well below their potentially optimal profits, even if they are in well balanced markets.
+Furthermore, these losses can increase significantly as markets become more imbalanced, getting to $20-40\%$ of optimal.
+Figure 4: Illustration of the gap from optimal profits that canonical A/B experimentation creates.
+8 An Empirical Case Study
+Thus far, we have shown that A/B pricing experiments can get the sign wrong, and that associated profit losses can be significant in a two-sided market model with logit demand.
+To align our analysis even more closely with practical scenarios, we empirically calibrate our model from the previous section using published data from several Airbnb markets, and investigate the conditions under which the sign-change phenomenon may occur in price experiments within this context.
+Although our model simplifies the complexities of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with real-world experiments.
+As will be shown in this section, our results suggest that a change of sign is indeed quite plausible in the empirical Airbnb context we study, demonstrating that these errors are not just a knife-edge issue—and thus potentially of high consequence for practitioners.
+8.1 Model Calibration
+In this section, we calibrate a specific version of our model from Section 7 to several Airbnb markets using the data from Farronato and Fradkin, (2022).
+Specifically, we consider the model from that section with $v(p)=e^{V-p}$.
+Note that this is a natural choice implied by a logit random utility model, where utility for the product is quasilinear in the price.
+For fixed $\varepsilon$, $V$, $\lambda$, and $p$, the resulting market balance equation is:
+$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)e^{V-p}}{\varepsilon+s^{*}(\lambda,p)e^{V-p}}.$ (28)
+Recall that then in this model, $D(p)=\rho-s^{*}(\lambda,p)$, and for a given supplier cost $c$, its profit is $\pi(p)=(p-c)D(p)$.
+In this exercise, we consider that $GTE$ corresponds to suppliers’ profits and that the experiment is on actual prices.
+It is worth noting that in practice, Airbnb does not actually set the price of listings, but instead sets fees.
+However, we have elected to use our price testing framework (in place of the fee-model detailed in Appendix 12) because the estimates made in Farronato and Fradkin, (2022) are better suited to examining price experiments.
+.
+Airbnb actively provides pricing guidance to hosts aimed at optimizing their pricing strategies and regularly conducts A/B tests to evaluate these recommendations (Deng and Le, 2023).
+This setting aligns closely with the price testing framework employed in this section.
+Furthermore, since fee changes are often passed on to consumers through adjustments to prices, this is a reasonable approach for our analysis.
+The goal of our calibration process is to arrive at imputed values of the parameters that define the model: (1) the total mass of listings $\rho$; (2) the outside option value $\varepsilon$; (3) the cost $c$; (4) the value $V$; (5) the arrival rate $\lambda$; and (6) the reference price $p$, from which the platform runs a price experiment.
+We now describe how we arrive at these six parameters.
+We leverage structural estimates made by Farronato and Fradkin, (2022) for Airbnb markets in major U.S.
+cities.
+In particular, Farronato and Fradkin, (2022) segment the Airbnb market into four tiers: economy, midscale, upscale, and luxury.
+They study each of these tiers across 10 cities.
+We refer to each pair of tier and city as a segment, and use superscript $s$ to denote a segment.
+For each segment $s$, Farronato and Fradkin, (2022) reports (1) an estimate of the average cost (per night booked), which we denote $\hat{c}^{(s)}$; and (2) an estimate of own price elasticity of demand, which we denote $\hat{e}_{p}^{(s)}$.
+[^11] These values come directly from tables E8 and E14 in the Appendix of Farronato and Fradkin, (2022).
+For our empirical study, we treat each segment as its own instantiation of the market model in (28).
+For example, if we are currently considering midscale listings in Detroit, then $v(p)$ will represent the average consumer valuation for midscale Detroit listings at price $p$.
+On its part, $\varepsilon$ will represent the value of the “outside” option relative to this segment, accounting for all other Airbnb segments in that city and the option of not purchasing anything within the Airbnb marketplace (e.g., booking an hotel).
+Our calibration proceeds according to the following steps:
   1. 1.
-    $R_{0}(f_{0},f_{1},\vec{p},q)$ and $R_{1}(f_{0},f_{1},\vec{p},q)$ are both differentiable over their domain.
+    We set the total mass of listings $\rho=1$. This allows us to interpret $s^{*}(\lambda,p)$ as the fraction of listings that are available.
   2. 2.
-    $R_{1}(f,f,\vec{p},q)/q=R_{0}(f,f,\vec{p},q)/(1-q)=R(f,\vec{p}).$
-Let $f_{1}^{n}$ be a sequence such that $\lim_{n\to\infty}f_{1}^{n}=f_{0}$. We then write the canonical estimator $\widehat{GTE}_{\Pi}$ as follows:
-$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\lim_{n\to\infty}\frac{\Pi_{1}(f_{0},f_{1}^{n},\vec{p},q)/q-\Pi_{0}(f_{0},f_{1}^{n},\vec{p},q)/(1-q)}{f_{1}^{n}-f_{0}}.$
-This is equivalent to:
-$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\Pi_{1y}(f_{0},f_{0},\vec{p},q)/q-\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q),$ (30)
-where $\Pi_{0y}$ and $\Pi_{1y}$ represent the derivative of the control and treatment platform profit functions with respect to the second argument respectively.
-We define the bias as $Bias_{\Pi}(f_{0},\vec{p},q)=GTE_{\Pi}(f_{0},\vec{p},q)-\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)$.
-Using Assumption 8.1 we get that $\Pi(f_{0},\vec{p},q)=\Pi_{1}(f_{0},f_{0},\vec{p},q)/q$.
-We can then use an analogous argument to our derivation of (11) to write bias as:
-$Bias_{\Pi}(f_{0},\vec{p},q)=\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+\Pi_{1x}(f_{0},f_{0},\vec{p},q)/q.$ (31)
-Using out revenue functions and equation (28) we can rewrite the preceding expressions to be more useful. First we rewrite bias as:
-$Bias_{\Pi}(f_{0},\vec{p},q)=f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+f_{0}R_{1x}(f_{0},f_{0},\vec{p},q)/q.$
-Finally, we rewrite $\widehat{GTE}_{\Pi}$ from equation (30) as follows:
-$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=R_{1}(f_{0},f_{0},\vec{p},q)/q+f_{0}R_{1y}(f_{0},f_{0},\vec{p},q)/q-f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q).$
-As before, there are two necessary and sufficient conditions for the naive estimator to have the wrong sign in a fee experiment.
-Note that as was true in the pricing problem, bias is always positive when firms test a small price increase.
-This means the general conditions for a sign flip to occur are $GTE(f_{0},\vec{p})\geq 0$, meaning the true treatment effect of a small price increase is non-negative.
-And $Bias(f_{0},\vec{p},q)\geq\widehat{GTE}(f_{0},\vec{p},q)$, meaning there is enough bias to cause the estimated treatment effect to be non-positive.
-These two conditions can be expressed as follows. First:
-[…]
-And second:
-[…]
+    As usual, we normalize the value of the outside option, $\varepsilon=1$.
+  3. 3.
+    We set the cost $c=\hat{c}^{(s)}$.
+  4. 4.
+    We choose a reference price $p$ at which we want to evaluate price experiments.
+Actual prices were not reported in Farronato and Fradkin, (2022); in lieu of this information, in our results we vary $p$ in a “reasonable” range, as we describe below.
+  5. 5.
+    We choose an arrival rate $\lambda$.
+Again, no proxy for $\lambda$ is reported in Farronato and Fradkin, (2022), and so in our results we also vary $\lambda$ in a “reasonable” range, again as we describe below.
+  6. 6.
+    Finally, given the chosen values of $\rho,\ \varepsilon,\ p$ and $\lambda$, we choose $V$ so that the demand elasticity $pD^{\prime}(p)/D(p)$ for $D(p)=\rho-s^{*}(\lambda,p)$, with $s^{*}(\lambda,p)$ defined in (28), is equal to $\hat{e}_{p}^{(s)}$ (the estimated elasticity from Farronato and Fradkin, (2022)).
+It is straightforward to check the price elasticity of demand in our demand model is monotonically increasing in $V$, which guarantees that the fitted value of $V$ is uniquely determined in this approach.
+Observe that at the conclusion of this procedure, for each value of $\lambda$ and $p$, we have an imputed demand model.
+We then use this demand model to determine whether a local CR or LR price experiment yields the wrong sign on the estimate of $GTE_{\pi}$, as discussed throughout the paper.
+The two “free parameters” in our calibration are the reference price $p$, and the arrival rate $\lambda$, so we vary them in our subsequent empirical analysis.
+We vary the reference price $p$ in a plausible range from the cost $c$ to a value above the monopoly optimal price implied by $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$.
+In our results we also vary $\lambda$ in a plausible range; in particular, we vary $\lambda$ so that in our fitted model we obtain plausible occupancy rates.
+The occupancy rate is the fraction of total supply that is booked in equilibrium, defined as $\zeta=1-s^{*}(\lambda,p)/\rho$.
+We vary $\lambda$ to ensure that we plot occupancy rates from 5% to 95%.
+As an example, one estimate from external market research on short term rental properties Airdna, (2024) estimates that the average occupancy rate across the entire Airbnb market is approximately 60%, with rates closer to 65% for major cities.
+This suggests the range of values of $\zeta$ in our results includes realistic values.
+In our results, we present our visualizations with the two axes of price $p$ and occupancy rate $\zeta$, the latter in place of the arrival rate $\lambda$.
+We numerically verified that in this model, each pair of original parameters ($\lambda,p$), yields a unique occupancy pair ($\zeta,p$).
+Further, we have verified that in all instances of our model, as $\lambda$ increases, the resulting $\zeta$ increases monotonically as well.
+We make this transformation because the occupancy rate is a more natural metric for interpretation than the arrival rate of customers.
+8.2 Results
+Figure 5: Illustrations of the change of sign region (purple) in LR-experiments (top row) and CR-experiments (bottom row), respectively, for economy listings (left column) and luxury listings (right column) in Boston.
+The vertical axis is the imputed occupancy rate, and the horizontal axis is the (baseline) control price around which an experiment is conducted.
+The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red.
+The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
+We plot prices from the cost estimate $\hat{c}^{(s)}$, to a fixed amount above the critical price $p_{c}^{(s)}$.
+Figure 5 shows the results of our model’s estimates for LR (top row) and CR (bottom row) price experiments in Boston.
+The left two plots show results for the economy tier, while the right two plots show results for the luxury tier.
+We have chosen to include Boston as an example city, but it is important to note that the trends observed in Figure 5, hold for all segments.
+(See Appendix 11 for these additional figures for all other segments.)
+As in our preceding visualizations, the shaded blue region in each plot is where Condition (a) holds in the fitted demand model, cf.
+(13), and the shaded red region is where Condition (b) holds, cf.
+(14).
+These regions overlap in the purple region where both conditions hold.
+These are the regions where the estimated treatment effect would have the wrong sign.
+As noted in the preceding section, our results are plotted with the reference price $p$ on the x-axis (varied in a reasonable range, as described in the previous subsection), and the imputed occupancy rate $\zeta$ on the y-axis.
+First, observe that in each figure, the right boundary of the change of sign region is a vertical line at a critical price $p_{c}^{(s)}$.
+To see why, observe that Condition (a) in (13) holds with equality at price $p_{c}^{(s)}$, with the cost $\hat{c}^{(s)}$ and elasticity $\hat{e}_{p}^{(s)}$ corresponding to the segment $s$.
+In particular, $p_{c}^{(s)}$ satisfies the first order condition for optimal monopoly pricing given $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$.
+Further, this identity is insensitive to the value of the occupancy rate $\zeta$ (or equivalently, the arrival rate $\lambda$).
+Given the monotonicity in Condition (a), at any higher (resp., lower) values of $p$ than $p_{c}^{(s)}$, Condition (a) is guaranteed to not hold (resp., is guaranteed to hold); and again, these facts are also insensitive to the value of $\zeta$.
+More importantly, observe that the figures make clear that changes of sign are prevalent in this model.
+Notably, across the parameter values we tested, LR experiments in particular seem especially vulnerable to changes of sign.
+This is consistent with the fact that the observed occupancy rates suggest a relatively demand-constrained market, at least as far as bias in price experiments is concerned.
+On the other hand, CR experiments exhibit a more nuanced behavior.
+For such experiments, a change of sign is observed in markets that are highly supply-constrained (e.g., with greater than 80-85% of listings being occupied), and when the reference price is near but still below the critical price $p_{c}^{(s)}$.
+It should be noted that although CR price experiments exhibit more favorable behavior in these plots than LR price experiments, LR price experiments are more common in industry.
+In part this is because platforms are often reticent to offer individual customers different prices for the same product (Phillips, 2021).
+As a consequence, one potential practical takeaway from our work is that firms may in fact have chosen an experimental design that is more commonly producing the wrong sign in their estimated treatment effects in practice, at least in the settings we have studied.
+Despite the stylized nature of the model, the prevalence of changes of sign across a wide range of parameters suggests this is a phenomenon of significant consequence for practitioners.
 9 Conclusion and Future Work
-In this paper we studied biases that arise due to interference in platform experiments, with a particular emphasis on profit or revenue metrics.
-Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue.
-We then show a novel method for debiasing canonical A/B experiments, by splitting half of subjects into treatment and half into control.
-In future work we aim to extend our debiasing result to allow for correcting canonical experiments beyond the equal split between control and treatment.
-We also aim to run empirical studies to get a better sense of the prevalence of changes of sign in practice.
+In this paper we studied the biases that arise due to interference in platform experiments, with a particular emphasis on profit and revenue metrics.
+Notably, we find that in certain settings, these biases can be so severe that they introduce a “change of sign” into the canonical estimators for profit, suggesting to firms that they should move prices (or fees) in the wrong direction.
+We then derive a novel method for debiasing canonical A/B experiments, by splitting units equally between treatment and control.
+We conclude by giving empirical evidence for the plausibility of estimators with the wrong sign in practical settings, arguing that firms should be more cautious when using canonical A/B pricing experiments.
+In future work we aim to extend our debiasing result, to allow for the correction of canonical experiments in cases where the treatment/control split is not 50/50.
+In addition to being more flexible, this may allow us to mitigate the sampling inefficiency of our current debiasing method.
+Currently, our method only uses half of the experimental data, namely, only the observed control demand, which may be a source of increased variance.
+Finally, we aim to run empirical studies to get a better sense of the true prevalence of changes of sign in practice.

### v5

@@
 We apply these results to a two-sided market model, and demonstrate how the “change of sign” regime depends on market factors such as the supply/demand imbalance, and the price markup.
-We conclude by calibrating our two-sided market model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners.
-[^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his critical assistance for our Airbnb empirical case study.
-We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS Revenue Management and Pricing Conferencea, the INFORMS Annual Meeting, the INFORMS Manufacturing and Service Operations Management Conference, and Airbnb for helpful feedback.
+We conclude by calibrating our revenue-management pricing model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners.
+[^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his significant assistance on our Airbnb empirical case study.
+We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS RMP Conference, the INFORMS Annual Meeting, the INFORMS MSOM Conference, and Airbnb for helpful feedback.
 Orrie Page was supported by a National Science Foundation Graduate Research Fellowship.
@@
 Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below).
-This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
+This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2025, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
 To obtain additional insight, in Section 7, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022).
@@
 We also investigate numerically how bias depends on both market balance and the baseline control price, and show how canonical A/B methods for price experimentation may result in significant profit underperformance.
-In Section 8, we conclude the paper by calibrating our two-sided market model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S.
+In Section 8, we conclude the paper by calibrating our revenue-management pricing model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S.
 cities.
-This exercise simplifies the Airbnb marketplace in many dimensions, but we believe it is useful for offering insights into the prevalence of estimators with the wrong sign in real-world markets.
-In our investigation, we find that such estimators are plausible in the Airbnb context.
-Notably, all markets in the study yielded estimators with the wrong sign when we simulated LR experiments, which would be a typical approach for testing pricing changes in practice.
-Although this analysis is based on a simplified market model, the results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners.
+This exercise is useful for offering insights into the plausibility of estimators with the wrong sign in real-world markets.
+In our investigation, we find that such estimators are quite plausible in the Airbnb context.
+Notably, all markets in the Farronato and Fradkin, (2022) study demonstrated a susceptibility to estimators with the wrong sign in LR experimentation, which would be the typical approach for testing pricing changes in practice.
+These results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners.
 We conclude in Section 9.
@@
 Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014, Fradkin, 2019, Liu et al., 2020), showing its importance in practice.
-Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
-A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023).
-Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation.
-More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects).
+Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2025, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2023, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
+A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023), Delarue and Kleanthis, (2025).
+Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation, or Delarue and Kleanthis, (2025), which proposes a method for reducing interference bias in matching markets.
+More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique to optimize supply-side payments via gradient descent in a centralized marketplace; while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects).
 Roemheld and Rao, (2024) uses an empirical study to suggest that false positives may be inflated due to interference in pricing experiments.
@@
 In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign.
+On its part, Li et al., (2025) study A/B experiments comparing recommendation algorithms and shows how symbiosis bias—arising from shared training data between treatment and control—can result in flipping the sign of the estimated global treatment effect.
 Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper—those where a continuous variable (such as price) is incrementally varied—as marginal policy changes, and develops a framework to study them.
@@
 Therefore, an experiment is conducted to determine the direction in which the decision variable should be adjusted in order to improve the performance metric (a form of policy improvement).
-We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control.
+We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction 1q1-q to control.
 We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated to each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
@@
 Notably, this notation abstracts away from the specific experimental design (whether CR or LR).
-We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
+[^3] At this level of abstraction, even other experimental designs like switchbacks could be analyzed.
+We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $0<q<1$.
 The canonical experimental designs at each side of the market have an associated natural “canonical” local estimator.
@@
 {assumption}
-$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $10<q<$ 1.
+$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $0<q<1$.
 The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric.
@@
 To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric).
-Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$.
+Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $0<\alpha<1$, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$.
 If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition.
@@
 $=T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\ .$ (3)
-The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.2.
+The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.
 The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
@@
 $Bias(x_{0},q)$ $:=GTE(x_{0})-\widehat{GTE}(x_{0},q)$ (4)
-$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .$ (5)
+$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .\$(5)
 Studying bias in this local regime proves fruitful, as it leads to the following simple characterization.
@@
 $T(x_{0})=T_{1}(x_{0},x_{0},q)/q\ .$
-By Assumption 3.2 and application of the chain rule we get:
+By Assumption 3 and application of the chain rule we get:
 $T^{\prime}(x_{0})=T_{1x}(x_{0},x_{0},q)/q+T_{1y}(x_{0},x_{0},q)/q\ .$
@@
 That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
-We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction $1-q$ to control.
+We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction 1q1-q to control.
 We denote by $p_{0}$ the baseline (control) price and by $p_{1}$ the treatment price.
@@
 In contrast, in a LR experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
-We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper.
-{assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
-In addition, $p_{0},p_{1}\geq c$.
+We make the following assumption that implies Assumption 3 and that is kept throughout the paper.
+{assumption} The functions $D_{0}(p_{0},p_{1},q)$ and $D_{1}(p_{0},p_{1},q)$ are continuously differentiable in their first two arguments $(p_{0},p_{1})$, for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and all $q\in(0,1)$.
+In addition, prices satisfy $p_{0},p_{1}\geq c$, where $c\geq 0$, and we assume $0<q<1$.
 Note that the assumption precludes the case of perfect substitution between control and treatment products.
@@
   1. 1.
-    $D_{0}(p_{0},p_{1},q)$ is non-decreasing in $p_{1}$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
+    $D_{0}(p_{0},p_{1},q)$ is non-decreasing in $p_{1}$, for all $p_{0}\in\mathbb{R}_{+}$ and $0<q<1$.
   2. 2.
-    $D_{1}(p_{0},p_{1},q)$ is non-decreasing in $p_{0}$, for all $p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
+    $D_{1}(p_{0},p_{1},q)$ is non-decreasing in $p_{0}$, for all $p_{1}\in\mathbb{R}_{+}$ and $0<q<1$.
   3. 3.
-    $D_{0}(p,p,q)=D_{1}(p,p,q)=D(p)$, for all $p\in\mathbb{R}_{+}$ and $10<q<$ 1.
+    $D_{0}(p,p,q)=D_{1}(p,p,q)=D(p)$, for all $p\in\mathbb{R}_{+}$ and $0<q<1$.
 The first two statements in Assumption 4.2 are natural as they imply that control and treatment products are (imperfect) substitutes.
@@
 $=(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$ (11)
-$=(p_{0}-c)Bias_{D}(p_{0},q)\ .$ (12)
+$=(p_{0}-c)Bias_{D}(p_{0},q)\ .\$(12)
 As discussed in Proposition 3.1, bias increases as interference effects between treatment and control groups increase.
@@
 Proposition 5.1
-We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
+We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $0<q<1$.
 This result follows similar cannibalization arguments developed in previous work (Johari et al., 2022).
@@
 By definition, $GTE_{\pi}(p_{0})\geq 0$ if and only if $\pi^{\prime}(p_{0})\geq 0$, which after some simple manipulation yields the following standard “markup” inequality:
-[…]
+$\mbox{Condition (a): \ \ }A(p_{0}):=\frac{p_{0}-c}{p_{0}}\leq-\frac{1}{e_{p}(p_{0})}\ ,$ (13)
 where $e_{p}(p_{0})=D^{\prime}(p_{0})p_{0}/D(p_{0})$ is the price elasticity of demand.
@@
 Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
-[…]
+$\mbox{Condition (b): \ \ }-\frac{1}{e_{p}(p_{0})}\leq\frac{(p_{0}-c)}{p_{0}}\frac{\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)}{D^{\prime}(p_{0})}:=B(p_{0},q)\ .$ (14)
 Note that condition (b) is a “modified markup” inequality. We have the following result.
@@
   3. 3.
-    $A(p_{0})\leq B(p_{0},q)$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
+    $A(p_{0})\leq B(p_{0},q)$, for all $p_{0}\in\mathbb{R}_{+}$ and $0<q<1$.
 Proof 5.3
@@
 Hence, we can re-write condition (b) as:
-$-\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
+$\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
 which can be understood as the experimental analogue of a markup inequality.
@@
 Viewing the fee as the variable of interest for the platform’s profit optimization, we obtain analogous results to those in this section, including the bias characterization, the possibility of a change of sign, and interpretation of the conditions for a change of sign in terms of appropriate derivatives.
-This extension suggests our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.
+This extension suggests that our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.
 6 Debiasing Canonical Estimators
@@
 The state of the system at time $t$ is given by $s_{t}$, which denotes the mass of available listings available at time $t$.
-[…]
+On the demand side, there is a continuum of customers arriving to the system at rate $\lambda$.
+When a customer arrives, they form a consideration set consisting of all available listings at that point in time.
+[^7] It is straightforward to add a process in which customers sample listings to consider with the same probability and independently.
+We omit this probability to simplify notation.
+Given the consideration set, customers decide whether or not to book a listing according to a multinomial logit choice model with an outside option of value $\varepsilon>0$.
+[^8] Note that by assumption we will constrain the value of the outside option $\varepsilon$ to be strictly positive.
+The value of booking a listing is given by the function $v(p)$.
+Under this specification, the probability that a customer who arrives in state $s$ books a listing is equal to:
 $\frac{sv(p)}{\varepsilon+sv(p)}.$
@@
   4. 4.
-    $-pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$
+    $pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$
 […]
@@
 We normalize by $\lambda$ instead of the GTE because the GTE for profits converges to 0 around the optimal price.
-As $\lambda$ increases, the market becomes more supply-constrained and bias for a LR experiment decreases as competition between listings dissipates.
+As $\lambda$ increases, the market becomes more supply-constrained and bias for an LR experiment decreases as competition between listings dissipates.
 Figure 2 shows the normalized bias of demand and profits in a CR experiment and this dependence on $\lambda$ is naturally reversed.
@@
 We recall the two conditions for a change of sign to occur:
-[…]
+$\text{Condition (a)}\quad GTE_{\pi}\geq 0;\qquad\text{Condition (b)}\quad Bias_{\pi}\geq GTE_{\pi}.$
 Interestingly, observe that the bias terms derived in Section 7.4 do not depend on the treatment fraction $q$ for either LR or CR experiments.
@@
 Figure 3: Illustration of the change of sign region (purple) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$.
-The vertical axis is the (log of) the arrival rate, or market balance, paramter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted.
+The vertical axis is the (log of) the arrival rate, or market balance, parameter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted.
 The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red.
@@
 Note that this price corresponds to the optimum suggested by repeated canonical A/B experiments.
-This is because, to the right of that price point, the expriment suggests decreasing the price, and to the left, to increase it.
+This is because, to the right of that price point, the experiment suggests decreasing the price, and to the left, increasing it.
 The difference between these two prices (the true optimal price and the experimentally optimized price) can correspond to significant losses in overall profit.
 In Figure 4, we see the percent of true optimal profit that a firm will achieve if it uses canonical LR (left) or CR (right) experimentation, for a range of market balance scenarios.
+Figure 4: Illustration of the gap from optimal profits that canonical LR, and CR A/B experimentation yields, with $v(p)=e^{V-p}$, $V=5$, $c=1$, $\rho=1$, and $\epsilon=1$.
 What we find is that the profit achieved from optimization via experimentation can be far below the optimal profit, with the impact worsening in highly demand-constrained markets (for LR experiments) or demand-constrained markets (for CR experiments).
@@
 Furthermore, these losses can increase significantly as markets become more imbalanced, getting to $20-40\%$ of optimal.
-Figure 4: Illustration of the gap from optimal profits that canonical A/B experimentation creates.
 8 An Empirical Case Study
-Thus far, we have shown that A/B pricing experiments can get the sign wrong, and that associated profit losses can be significant in a two-sided market model with logit demand.
-To align our analysis even more closely with practical scenarios, we empirically calibrate our model from the previous section using published data from several Airbnb markets, and investigate the conditions under which the sign-change phenomenon may occur in price experiments within this context.
-Although our model simplifies the complexities of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with real-world experiments.
-As will be shown in this section, our results suggest that a change of sign is indeed quite plausible in the empirical Airbnb context we study, demonstrating that these errors are not just a knife-edge issue—and thus potentially of high consequence for practitioners.
-8.1 Model Calibration
-In this section, we calibrate a specific version of our model from Section 7 to several Airbnb markets using the data from Farronato and Fradkin, (2022).
-Specifically, we consider the model from that section with $v(p)=e^{V-p}$.
-Note that this is a natural choice implied by a logit random utility model, where utility for the product is quasilinear in the price.
-For fixed $\varepsilon$, $V$, $\lambda$, and $p$, the resulting market balance equation is:
-$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)e^{V-p}}{\varepsilon+s^{*}(\lambda,p)e^{V-p}}.$ (28)
-Recall that then in this model, $D(p)=\rho-s^{*}(\lambda,p)$, and for a given supplier cost $c$, its profit is $\pi(p)=(p-c)D(p)$.
-In this exercise, we consider that $GTE$ corresponds to suppliers’ profits and that the experiment is on actual prices.
-It is worth noting that in practice, Airbnb does not actually set the price of listings, but instead sets fees.
-However, we have elected to use our price testing framework (in place of the fee-model detailed in Appendix 12) because the estimates made in Farronato and Fradkin, (2022) are better suited to examining price experiments.
-.
-Airbnb actively provides pricing guidance to hosts aimed at optimizing their pricing strategies and regularly conducts A/B tests to evaluate these recommendations (Deng and Le, 2023).
-This setting aligns closely with the price testing framework employed in this section.
-Furthermore, since fee changes are often passed on to consumers through adjustments to prices, this is a reasonable approach for our analysis.
-The goal of our calibration process is to arrive at imputed values of the parameters that define the model: (1) the total mass of listings $\rho$; (2) the outside option value $\varepsilon$; (3) the cost $c$; (4) the value $V$; (5) the arrival rate $\lambda$; and (6) the reference price $p$, from which the platform runs a price experiment.
-We now describe how we arrive at these six parameters.
-We leverage structural estimates made by Farronato and Fradkin, (2022) for Airbnb markets in major U.S.
-cities.
-In particular, Farronato and Fradkin, (2022) segment the Airbnb market into four tiers: economy, midscale, upscale, and luxury.
-They study each of these tiers across 10 cities.
-We refer to each pair of tier and city as a segment, and use superscript $s$ to denote a segment.
-For each segment $s$, Farronato and Fradkin, (2022) reports (1) an estimate of the average cost (per night booked), which we denote $\hat{c}^{(s)}$; and (2) an estimate of own price elasticity of demand, which we denote $\hat{e}_{p}^{(s)}$.
-[^11] These values come directly from tables E8 and E14 in the Appendix of Farronato and Fradkin, (2022).
-For our empirical study, we treat each segment as its own instantiation of the market model in (28).
-For example, if we are currently considering midscale listings in Detroit, then $v(p)$ will represent the average consumer valuation for midscale Detroit listings at price $p$.
-On its part, $\varepsilon$ will represent the value of the “outside” option relative to this segment, accounting for all other Airbnb segments in that city and the option of not purchasing anything within the Airbnb marketplace (e.g., booking an hotel).
-Our calibration proceeds according to the following steps:
-  1. 1.
-    We set the total mass of listings $\rho=1$. This allows us to interpret $s^{*}(\lambda,p)$ as the fraction of listings that are available.
-  2. 2.
-    As usual, we normalize the value of the outside option, $\varepsilon=1$.
-  3. 3.
-    We set the cost $c=\hat{c}^{(s)}$.
-  4. 4.
-    We choose a reference price $p$ at which we want to evaluate price experiments.
-Actual prices were not reported in Farronato and Fradkin, (2022); in lieu of this information, in our results we vary $p$ in a “reasonable” range, as we describe below.
-  5. 5.
-    We choose an arrival rate $\lambda$.
-Again, no proxy for $\lambda$ is reported in Farronato and Fradkin, (2022), and so in our results we also vary $\lambda$ in a “reasonable” range, again as we describe below.
-  6. 6.
-    Finally, given the chosen values of $\rho,\ \varepsilon,\ p$ and $\lambda$, we choose $V$ so that the demand elasticity $pD^{\prime}(p)/D(p)$ for $D(p)=\rho-s^{*}(\lambda,p)$, with $s^{*}(\lambda,p)$ defined in (28), is equal to $\hat{e}_{p}^{(s)}$ (the estimated elasticity from Farronato and Fradkin, (2022)).
-It is straightforward to check the price elasticity of demand in our demand model is monotonically increasing in $V$, which guarantees that the fitted value of $V$ is uniquely determined in this approach.
-Observe that at the conclusion of this procedure, for each value of $\lambda$ and $p$, we have an imputed demand model.
-We then use this demand model to determine whether a local CR or LR price experiment yields the wrong sign on the estimate of $GTE_{\pi}$, as discussed throughout the paper.
-The two “free parameters” in our calibration are the reference price $p$, and the arrival rate $\lambda$, so we vary them in our subsequent empirical analysis.
-We vary the reference price $p$ in a plausible range from the cost $c$ to a value above the monopoly optimal price implied by $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$.
-In our results we also vary $\lambda$ in a plausible range; in particular, we vary $\lambda$ so that in our fitted model we obtain plausible occupancy rates.
-The occupancy rate is the fraction of total supply that is booked in equilibrium, defined as $\zeta=1-s^{*}(\lambda,p)/\rho$.
-We vary $\lambda$ to ensure that we plot occupancy rates from 5% to 95%.
-As an example, one estimate from external market research on short term rental properties Airdna, (2024) estimates that the average occupancy rate across the entire Airbnb market is approximately 60%, with rates closer to 65% for major cities.
-This suggests the range of values of $\zeta$ in our results includes realistic values.
-In our results, we present our visualizations with the two axes of price $p$ and occupancy rate $\zeta$, the latter in place of the arrival rate $\lambda$.
-We numerically verified that in this model, each pair of original parameters ($\lambda,p$), yields a unique occupancy pair ($\zeta,p$).
-Further, we have verified that in all instances of our model, as $\lambda$ increases, the resulting $\zeta$ increases monotonically as well.
-We make this transformation because the occupancy rate is a more natural metric for interpretation than the arrival rate of customers.
-8.2 Results
-Figure 5: Illustrations of the change of sign region (purple) in LR-experiments (top row) and CR-experiments (bottom row), respectively, for economy listings (left column) and luxury listings (right column) in Boston.
-The vertical axis is the imputed occupancy rate, and the horizontal axis is the (baseline) control price around which an experiment is conducted.
-The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red.
-The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
-We plot prices from the cost estimate $\hat{c}^{(s)}$, to a fixed amount above the critical price $p_{c}^{(s)}$.
-Figure 5 shows the results of our model’s estimates for LR (top row) and CR (bottom row) price experiments in Boston.
-The left two plots show results for the economy tier, while the right two plots show results for the luxury tier.
-We have chosen to include Boston as an example city, but it is important to note that the trends observed in Figure 5, hold for all segments.
-(See Appendix 11 for these additional figures for all other segments.)
-As in our preceding visualizations, the shaded blue region in each plot is where Condition (a) holds in the fitted demand model, cf.
-(13), and the shaded red region is where Condition (b) holds, cf.
-(14).
-These regions overlap in the purple region where both conditions hold.
-These are the regions where the estimated treatment effect would have the wrong sign.
-As noted in the preceding section, our results are plotted with the reference price $p$ on the x-axis (varied in a reasonable range, as described in the previous subsection), and the imputed occupancy rate $\zeta$ on the y-axis.
-First, observe that in each figure, the right boundary of the change of sign region is a vertical line at a critical price $p_{c}^{(s)}$.
-To see why, observe that Condition (a) in (13) holds with equality at price $p_{c}^{(s)}$, with the cost $\hat{c}^{(s)}$ and elasticity $\hat{e}_{p}^{(s)}$ corresponding to the segment $s$.
-In particular, $p_{c}^{(s)}$ satisfies the first order condition for optimal monopoly pricing given $\hat{c}^{(s)}$ and $\hat{e}_{p}^{(s)}$.
-Further, this identity is insensitive to the value of the occupancy rate $\zeta$ (or equivalently, the arrival rate $\lambda$).
-Given the monotonicity in Condition (a), at any higher (resp., lower) values of $p$ than $p_{c}^{(s)}$, Condition (a) is guaranteed to not hold (resp., is guaranteed to hold); and again, these facts are also insensitive to the value of $\zeta$.
-More importantly, observe that the figures make clear that changes of sign are prevalent in this model.
-Notably, across the parameter values we tested, LR experiments in particular seem especially vulnerable to changes of sign.
-This is consistent with the fact that the observed occupancy rates suggest a relatively demand-constrained market, at least as far as bias in price experiments is concerned.
-On the other hand, CR experiments exhibit a more nuanced behavior.
-For such experiments, a change of sign is observed in markets that are highly supply-constrained (e.g., with greater than 80-85% of listings being occupied), and when the reference price is near but still below the critical price $p_{c}^{(s)}$.
-It should be noted that although CR price experiments exhibit more favorable behavior in these plots than LR price experiments, LR price experiments are more common in industry.
-In part this is because platforms are often reticent to offer individual customers different prices for the same product (Phillips, 2021).
-As a consequence, one potential practical takeaway from our work is that firms may in fact have chosen an experimental design that is more commonly producing the wrong sign in their estimated treatment effects in practice, at least in the settings we have studied.
-Despite the stylized nature of the model, the prevalence of changes of sign across a wide range of parameters suggests this is a phenomenon of significant consequence for practitioners.
+In this section, we investigate the feasibility of changes of sign in price experments in a calibrated, real-world setting.
+In particular, we consider the lodging marketplace Airbnb, using market conditions for a number of cities and market segments estimated by Farronato and Fradkin, (2022).
+We leverage their estimates to calibrate our change of sign conditions (13) and (14).
+As we show, the estimates of Farronato and Fradkin, (2022) allow us to obtain interpretable conditions under which a change of sign might occur in a price experiment in each city and market segment.
+Notably, we demonstrate that for a reasonable range of market conditions, changes of sign would be expected in practice.
+We begin by rewriting our change of sign conditions to be more suited to the Farronato and Fradkin, (2022) data that estimates own price elasticities for (city, market) segments.
+Recall the following two necessary and sufficient conditions for a change of sign to occur for the canonical price experiment we study, cf.
+(13)-(14):
+Condition (a): $GTE_{\pi}(p)\geq 0;$
+Condition (b): $Bias_{\pi}(p,q)\geq GTE(p).$
+Note that $e_{p}(p)$ is the price elasticity of demand and is defined as $e_{p}(p):=pD^{\prime}(p)/D(p)$.
+The first of these conditions, Condition (a), ensures that the true global treatment effect of a sufficiently small price increase is positive.
+In other words, this condition is only satisfied when the firm is currently (locally) “underpricing” relative to the optimal price .
+We can rewrite this condition using the monopoly pricing optimality condition as follows:
+$GTE_{\pi}(p)\geq 0\iff$ $\frac{p}{p-c}+e_{p}(p)\geq 0.$ (28)
+The second of these conditions, Condition (b), is satisfied when the interference bias is at least as large as the GTE itself.
+For the purposes of this exercise, we suppose that $q=1/2$, i.e., that half of all units are treated, and half are left in control.
+As discussed in Section 6, under this equal allocation experimental design, there holds $D_{1x}(p,p,q)=D_{0y}(p,p,q)$.
+In order to rewrite Condition (b), we take advantage of (11):
+$Bias_{\pi}(p,q)=(p-c)\left(D_{0y}(p,p,q)+D_{1x}(p,p,q)\right).$
+Combining this with the assumption that $q=1/2$, we can rewrite bias as:
+$Bias_{\pi}(p,1/2)=2(p-c)D_{0y}(p,p,1/2).$
+To proceed, define $e_{0y}(p,p,q)$ as the cross-price elasticity of control demand, i.e., the percentage change in control demand for a one percent increase in the treatment price.
+Recalling that $D_{0}(p,p,q)=Q_{0}(p,p,q)/(1-q)$, we define $e_{0y}(p,p,q)$ as follows:
+$e_{0y}(p,p,q):=\frac{Q_{0y}(p,p,q)p}{Q_{0}(p,p,q)}=\frac{D_{0y}(p,p,q)p}{D_{0}(p,p,q)}.$
+We can then write Condition (b) as:
+$Bias_{\pi}(p,1/2)\geq GTE_{\pi}(p)\iff$ $2(p-c)D_{0y}(p,p,1/2)\geq D(p)+(p-c)D^{\prime}(p)$
+$\iff$ $\frac{2(p-c)}{p}\frac{pD_{0y}(p,p,1/2)}{D(p)}\geq 1+\frac{(p-c)}{p}\frac{pD^{\prime}(p)}{D(p)}$
+$\iff$ $\frac{2(p-c)}{p}e_{0y}(p,p,1/2)\geq 1+\frac{(p-c)}{p}e_{p}(p)$
+$\iff$ $2e_{0y}(p,p,1/2)\geq\frac{p}{p-c}+e_{p}(p)$ (29)
+Note that $D_{0}(p,p,q)=D(p)$ by Assumption 4.2. Now, combining (28)-(29) allows us to write the change of sign conditions as the following chain of inequalities:
+$2e_{0y}(p,p,1/2)\geq\frac{p}{p-c}+e_{p}(p)\geq 0$ (30)
+Next, we investigate the feasibility of this chain of inequalities holding in real world markets.
+Observe that in (30) there are four unknown variables: the price $p$, the cost $c$, the own price elasticity of control demand $e_{p}(p)$, and the cross-price elasticity of demand $e_{0y}(p,p,1/2)$.
+Farronato and Fradkin, (2022) reports estimates of the costs and own price elasticities of demand for Airbnb market segments (economy, midscale, upscale, luxury) in 10 cities around the United States.
+This leaves us with two missing variables: the price and the cross-price elasticity.
+It is important to note that the cross-price elasticity, $e_{0y}(p,p,1/2)$, has a different interpretations in a two-sided market like Airbnb when running LR and CR experiments.
+In an LR experiment, $e_{0y}(p,p,1/2)$ quantifies the change in demand for control listings when the price of treated listings is raised.
+The experiment effectively creates two categories of products with higher (treatment) and lower (control) prices.
+If treatment raises the price, then control demand rises because price sensitive customers move away from the higher priced treatment listings, towards the control listings.
+The cross-price elasticity captures this effect, exactly as usual in demand modeling.
+In a CR experiment, $e_{0y}(p,p,1/2)$ has a different interpretation.
+Now, $e_{0y}(p,p,1/2)$ quantifies the degree to which demand from control customers changes when the price is raised for treatment customers.
+The mechanism by which this occurs depends on an availability effect: if the treatment raises the price, then as demand from treatment customers falls, product availability increases, which in turn positively impacts control demand.
+To demonstrate this more formally, with a slight abuse of notation, let $D(p,s)$ represent demand when price is $p$, and available inventory is $s$ (out of total inventory $\rho$).
+If we run a CR experiment on a fraction $q$ of customers, the following balance equation must hold:
+$(1-q)D(p_{0},s)+qD(p_{1},s)=\rho-s$
+This equation implicitly defines $s$ in terms of $p_{0},p_{1}$, and $q$.
+Indeed, observe that this a more general version of the CR steady-state condition given for the logit demand model in (25).
+If we now differentiate with respect to $p_{1}$, we obtain:
+$(1-q)D_{s}(p_{0},s)\frac{\partial s}{\partial p_{1}}+q\left[D_{p}(p_{1},s)+D_{s}(p_{1},s)\frac{\partial s}{\partial p_{1}}\right]=-\frac{\partial s}{\partial p_{1}}$
+Setting $p_{1}=p_{0}$ gives:
+$D_{s}(p_{0},s)\frac{\partial s}{\partial p_{1}}=-\frac{\partial s}{\partial p_{1}}-qD_{p}(p_{0},s)$ (31)
+In (31), the left hand side represents the change in rescaled control demand as the treatment price is raised; this is $D_{0y}(p,p,q)$ for CR experiments.
+The right hand side shows that this quantity depends on two effects: a direct price effect on demand, $D_{p}(p_{0},s)$; but also the impact of the price change on equilibrium availability, $\partial s/\partial p_{1}$.
+Measuring this quantity requires empirical calibration of the change in supply rationing when prices are changed for a subset of customers.
+With published estimates such as those in Farronato and Fradkin, (2022), it is difficult to estimate any reasonable range for this availability effect.
+We therefore focus on LR experimentation in our presentation, rather than CR experimentation.
+Although $e_{0y}(p,p,1/2)$ is not observed, it is possible to consider a reasonable range for this cross elasticity, by comparing it to the underlying own price elasticity of demand in each market segment as reported in Farronato and Fradkin, (2022).
+As we have discussed above, LR price experimentation is the industry standard.
+Therefore this approach also captures the regime of most interest to practitioners.
+Returning now to the chain of inequalities which produce a sign flip (30), we can use the data reported in Farronato and Fradkin, (2022) to plot ranges of prices and cross-price elasticities for which we would expect to see changes of sign in canonical LR price experiments.
+For price, we consider a range beginning at cost, and ending slightly above the optimal monopoly price implied by $\hat{e}_{p}$ and $\hat{c}$.
+For the cross-price elasticity we plot beginning at 0, and end at $2\times-\hat{e}_{p}$.
+We choose this range of cross-price elasticities because in general we expect the cross-price elasticity to be commensurate with the magnitude of the own price elasticity in LR experiments.
+Since the true cross-price elasticity is unknowable given our data, we select this range to be conservative.
+We should note that while it is possible to have a cross-price elasticity that exceeds $2\times-\hat{e}_{p}$, in these settings the interference bias will be so large that a sign flip is even more likely if the firm is underpricing.
+Figure 5: Change of sign region for price and cross-price elasticity in the Airbnb Boston Economy Segment.
+Farronato and Fradkin, (2022) estimates $\hat{c}=\mathdollar 81.4$, and $\hat{e}_{p}(p):-2.33$.
+Figure 5 shows the change of sign region plot for the economy segment of Airbnb listings in Boston.
+(See Appendix 11 for similar plots for the other market segments, as well as the other cities we consider; the results are similar to Figure 5.) In the blue region the bias is too low to cause a change of sign (Condition (b) is violated), while in the red region the current price is already too high (Condition (a) is violated).
+For parameters in the purple region, our analysis suggests that changes of sign will occur.
+It is important to note that since we do not observe the market price or the cross-price elasticity of demand, the estimates in Farronato and Fradkin, (2022) are insufficient to identify the exact state of the world on this plot.
+Nevertheless, one can reasonably conclude from Figure 5 that changes of sign are plausible.
+In particular, if listings are even mildly underpriced, and the cross-price elasticity of control demand is comparable to the magnitude of the own-price elasticity of demand (i.e., approximately $2$), then our results show that changes of sign would arise: the experiment will incorrectly suggest that the current price is too high, and the experimenter will be encouraged to move the price in the wrong direction.
+We view these results as compelling evidence that changes of sign are a phenomenon that can have significant consequence for practitioners.
 9 Conclusion and Future Work
-In this paper we studied the biases that arise due to interference in platform experiments, with a particular emphasis on profit and revenue metrics.
+In this paper we studied the biases that arise due to interference in platform pricing experiments, with a particular emphasis on profit and revenue metrics.
 Notably, we find that in certain settings, these biases can be so severe that they introduce a “change of sign” into the canonical estimators for profit, suggesting to firms that they should move prices (or fees) in the wrong direction.
@@
 We conclude by giving empirical evidence for the plausibility of estimators with the wrong sign in practical settings, arguing that firms should be more cautious when using canonical A/B pricing experiments.
-In future work we aim to extend our debiasing result, to allow for the correction of canonical experiments in cases where the treatment/control split is not 50/50.
-In addition to being more flexible, this may allow us to mitigate the sampling inefficiency of our current debiasing method.
-Currently, our method only uses half of the experimental data, namely, only the observed control demand, which may be a source of increased variance.
-Finally, we aim to run empirical studies to get a better sense of the true prevalence of changes of sign in practice.
+We propose multiple direction for future work.
+First, we believe that more analysis needs to be done on the debiasing of canonical estimators.
+Our debiasing method, while novel, relies on the treatment/control split being precisely 50/50.
+A debiasing method that functions in all treatment split settings would be both more flexible, and would allow us to mitigate the sampling inefficiency of our current debiasing method.
+Our current debiasing approach only uses half of the available data, namely only the control demand, which may be a source of increased variance in practice.
+A second, and related direction for future work involves quantifying the variance of these estimators themselves.
+As we note in Section 7, an interval of prices always exists in which a sign flip will occur in expectation (for a logit demand function), but this interval is bordered by the optimal price.
+In close-to-optimal settings, it is possible for estimators to be assigned the wrong sign entirely by fault of natural variance in the experiment.
+Typically, practitioners avoid this by dismissing results that are not statistically significant, meaning many of these incorrectly signed estimators may be ignored in practice because they fail to reach some critical p-value.
+Identifying conditions under which we would expect to see a statistically significant estimator with the wrong sign would be helpful to fully quantify how these sign flips actually impact practitioners.
+That said, our view is that practitioners should be more concerned with the impact of bias as studied in our paper, since variance vanishes with large sample sizes while bias does not.
+Finally, our paper motivates the need for further empirical study into this phenomenon.
+Our empirical results suggest that estimators with the wrong sign may be more prevalent than previously expected.
+One approach might be for practitioners to employ a ”meta-experiment” similar to Holtz et al., (2025), with both a canonical LR experiment and a cluster-randomized experiment, as a way to compare the usual approach to a relatively less biased alternative.
+More generally, our work suggests there is benefit in developing empirical tests to identify when debiasing is needed to avoid making wrong decisions.